### v1

# Conditioned stochastic stability of equilibrium states on uniformly hyperbolic sets

Abstract.
We establish conditioned stochastic stability of equilibrium states for Hölder weights on uniformly hyperbolic sets.
Key words and phrases:
Stochastic stability; absorbing Markov processes; conditioned random dynamics; thermodynamic formalism; transient dynamics
2020 Mathematics Subject Classification:
60J05; 37C30; 37D20; 37D35; 37D45
1. Introduction
The long-term statistical behaviour of trajectories for typical initial conditions in a dynamical system is well-known to be characterised by ergodic measures. Indeed, given a map $T:M\to M$ and a $T$-invariant ergodic measure $\nu$, Birkhoff’s ergodic theorem states that, for any bounded measurable observables $\varphi:M\to\mathbb{R}$, “time average equal space average” in the sense that
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\varphi\circ T^{i}(x)=\int\varphi\mathrm{d}\nu,\quad\nu\text{-a.s. on }x.$ (1)
Importantly, “typical initial conditions” refers to (1) being a “$\nu$-almost sure” statement and highlights the possible coexistence of many (even uncountably many) ergodic invariant measures. In this context, the notion of stochastic stability provides a strategy for identifying those ergodic invariant measures that persist under small random perturbations. Stochastic stability consists of (i) proving the existence and uniqueness of a stationary measure $\eta_{\varepsilon}$ for the process $X_{\varepsilon}$ generated by $\varepsilon$-small random perturbations of $T$, and (ii) characterising the limit (in weak-) of $\eta_{\varepsilon}$ as $\varepsilon\to 0$. If $\eta_{\varepsilon}\to\eta$ as $\varepsilon\to 0$, then $\eta$ is the only ergodic invariant measure of $T$ robust under small random perturbation and we may say that $\eta$ is stochastically stable. By construction, this strategy is fit to identify invariant measures sitting on attractors, near which $\eta_{\varepsilon}$ accumulates most mass. [^1] This is not always the case, see for example [Gora1985] or [AraujoTahzibi2005] where the stochastically stable measure of LSV maps with $\alpha\geq 1$ is $\delta_{0}$, which sits on a non-uniformly hyperbolic repeller.
Repelling invariant sets may also support many ergodic invariant measures. Nevertheless, the stochastic stability of such measures cannot be characterised with the strategy presented above since the stationary measure of $X_{\varepsilon}$ will often attribute little-to-no mass near such sets and instead will highlight attractors. To address this issue, we introduced the notion of “conditioned stochastic stability” in [BassolsCastroLamb2024] where, instead of stationary measures, we proposed considering quasi-ergodic measures of the process $X_{\varepsilon}$ conditioned upon remaining near the repeller.
In essence, conditioned stochastic stability consists of: (i) proving the existence and uniqueness of a quasi-ergodic measure $\nu_{\varepsilon}$ for the process $X_{\varepsilon}$ generated by $\varepsilon$-small random perturbations of $T$ and conditioned upon remaining in a neighbourhood of a repelling set $\Lambda$, and (ii) characterising the limit (in weak-) of $\nu_{\varepsilon}$ as $\varepsilon\to 0$. In [BassolsCastroLamb2024], we applied this strategy to uniformly expanding sets $\Lambda$ and identified that quasi-ergodic measures converge to so-called equilibrium states in the limit of $\varepsilon\to 0$, thus being conditionally stochastically stable
In this paper, we extend our previous results from [BassolsCastroLamb2024] to the uniformly hyperbolic setting, allowing for a contracting direction. While this is a natural generalisation of [BassolsCastroLamb2024], it is a non-trivial task to adapt the existing results to the hyperbolic setting. In particular, the dynamics do not allow for standard Hilbert cone techniques to obtain quasi-ergodic measures nor control their support, thus hindering the proof of the existence of quasi-ergodic measures. This is also a consequence of the transitivity of $T$ not assumed to hold in a neighbourhood of any basic set. To address these issues, here we resort to spectral techniques on the anisotropic Banach spaces $B^{t,s}$ of Baladi and Tsujii [Baladi-Tsujii2008, Baladi2018-book], together with perturbative results from Keller and Liverani [KellerLiverani1999]. Moreover, we provide a more detailed analysis of the global conditioned dynamics and global quasi-ergodic measures for systems with multiple repellers (see the discussion in Section 2.3 and Theorem 6.5 in particular).
1.1. Conditioned Markov processes
Given a Markov chain on a metric space $(E,d)$, suppose that we are only interested in its behaviour as long as it remains in a compact subset $M\subset E$. We thus identify $E\setminus M$ with a cemetery state $\partial$ and consider the state space $E_{M}\coloneqq M\sqcup\partial$ with the induced topology. On $E_{M}$, we consider the absorbing Markov chain absorbed in $\partial$ and given by
$X\coloneqq(\Omega,\{\mathcal{F}_{n}\}_{n\in\mathbb{N}_{0}},\{X_{n}\}_{n\in\mathbb{N}_{0}},\{\mathbf{P}^{n}\}_{n\in\mathbb{N}_{0}},\{\mathbb{P}_{x}\}_{x\in E_{M}})$
in the usual sense (see e.g [Rogers1994, Definition III.1.1]). Since $\partial$ is a cemetery state, we impose $\mathbf{P}^{n}(\partial,\partial)=1$ for all $n\in\mathbb{N}_{0}$, i.e. $X_{n}$ is absorbed in $\partial$.
Given a non-positive continuous potential function $\phi:M\to\mathbb{(}-\infty,0]$, we lift $X$ to the new process $X^{\phi}$ given by (see also [Gouezel-Liverani2008, Section 8.1.2] and [Baladi2018-book, Remark 5.23])
$X_{n+1}^{\phi}=\begin{cases}X_{n+1},&\text{with prob. }e^{\phi(X_{n})}\\ \partial,&\text{with prob. }1-e^{\phi(X_{n})}.\end{cases}$
The process $X^{\phi}$ is again a Markov chain, now with transition kernels $\mathbf{P}_{\phi}(x,\cdot)\coloneqq e^{\phi(x)}\mathbf{P}(x,\cdot)$ and inheriting the naturally induced family of measures $\mathbb{P}_{x}^{\phi}$ on $E_{M}$ from $X$. Observe that $X_{n}=X_{n}^{\phi}$ and $\mathbb{P}_{x}=\mathbb{P}_{x}^{\phi}$ if $\phi(x)=0$ for every $x\in M.$ We denote by $\mathbb{E}_{x}$ and $\mathbb{E}_{x}^{\phi}$ the expectation of the process $X_{n}$ and $X_{n}^{\phi}$ with respect to the probability measures $\mathbb{P}_{x}$ and $\mathbb{P}_{x}^{\phi}$, respectively. Here, we are interested in the behaviour of $X_{n}^{\phi}$ before it is absorbed by $\partial$ or, in other words, before the stopping time
$\tau(\omega,x)\coloneqq\inf\{n>0:\,X_{n}^{\phi}\in\partial\},\quad(\omega,x)\in\Omega\times M$
occurs. As mentioned above, quasi-ergodic measures are defined to capture the statistical behaviour of the process $X_{n}^{\phi}$ conditioned upon not being absorbed, i.e. conditioned upon $\tau>n$. To be more specific:
Definition 1.1 (Quasi-ergodic measure on $\mathcal{U}\subset M$).
We say that the probability measure $\nu$ on $M$ is a quasi-ergodic measure on $\mathcal{U}\subset M$ for the $e^{\phi}$-weighted process $X^{\phi}$ if for any observable $\varphi:\mathcal{U}\to\mathbb{R}$ it holds that
$\begin{split}\mathbb{E}_{x}^{\phi}\left[\frac{1}{n}\sum_{i=0}^{n-1}\varphi\circ X_{i}^{\phi}\,\bigg{|}\,\tau>n\right]\xrightarrow[]{n\to\infty}\int\varphi\mathrm{d}\nu,\quad\nu\text{-a.s. on }x,\end{split}$ (2)
where $\partial\coloneqq M\setminus\mathcal{U}$, i.e. $\tau(x,\omega)\coloneqq\inf\{n>0:\,X_{n}^{\phi}\not\in\mathcal{U}\},$ with the conditional expectation on the left-hand side defined as
$\frac{1}{\mathbb{P}_{x}^{\phi}\left[\tau>n\right]}\mathbb{E}_{x}^{\phi}\left[\mathbbm{1}_{\{\tau>n\}}\frac{1}{n}\sum_{i=0}^{n-1}\varphi\circ X_{i}^{\phi}\right]=\frac{1}{\mathbb{E}_{x}[e^{S_{n}\phi}\mathbbm{1}_{M}]}\mathbb{E}_{x}\left[e^{S_{n}\phi}\mathbbm{1}_{M}(X_{n})\frac{1}{n}\sum_{i=0}^{n-1}\varphi\circ X_{i}\right],$ (3)
and where $S_{n}\phi=\sum_{i=0}^{n-1}\phi\circ X_{i}$ denotes the Birkhoff sum.
Observe that this definition is well posed for any continuous $\phi$ and not only non-positive weights. For a general $\phi$, we interpret $X^{\phi}$ as follows: the process evolves in time and carries a certain amount of mass. If the process enters (dynamically) the region $\partial$, then it is killed and all mass is lost. At every time step $n$ and in every position $X_{n}$, the process will increase (or decrease) its mass by a factor $e^{\phi(X_{n})}$. Conditioning upon $\tau>n$ as in (2) must then be interpreted as the right-hand side in (3).
The study of Markov processes conditioned upon never entering a cemetery state has long been present in the literature [Yaglom1947, Pinsky1985, Breyer1999, Collet2013] and more recently become of increasing interest [Champagnat2016, Champagnat2018, Colonius2021, Castro2024, Castro2022]. To prove the existence and the uniqueness of the quasi-ergodic measure for a process $X_{\varepsilon}$, it is often required for the chain to be strong Feller and transitive (see e.g. [BassolsCastroLamb2024, Appendix A]). Nevertheless, this assumption is no longer valid when considering small random perturbations of hyperbolic (non-expanding) deterministic maps, as done herein. To address this shortcoming, we provide new results guaranteeing the existence and uniqueness of quasi-ergodic measures in the small-noise regime, whose support contains relevant dynamical information (see Theorems 2.6 and 2.8 guaranteeing $\Lambda\subset\mathrm{supp}\,\nu$).
1.2. Conditioned stochastic stability
The notion of conditioned stochastic stability presented in [BassolsCastroLamb2024] provides a natural candidate to highlight the statistics of the transient behaviour of a map $T$. As mentioned above, we may consider the quasi-ergodic measure of the Markov process $X_{\varepsilon}$ generated by small, $\varepsilon$-bounded, random perturbations of a map $T:M\to M$ and conditioned upon remaining in a particular region of interest, e.g. where the transient evolves. As $\varepsilon\to 0$, if the ($-^{*}$) limit [^2] While we only consider weak- convergence, one may also ask for the convergence in total variation of the quasi-ergodic measure, which would refer to strong conditioned stochastic stability [Shen-vanStrien2013]. of this quasi-ergodic measure exists and is unique, we have effectively identified a $T$-invariant measure [Khasminskii1963] that is robust under small random perturbations.
We distinguish the complement of two regions in state space with the potential to highlight dynamical transients: (i) $\partial=E\setminus\overline{V}$, where $V$ is the (isolating) neighbourhood of an invariant set (see Definition 2.2), and (ii) $\partial=\mathcal{A},$ an open neighbourhood of a forward invariant set for $T$ which we may think of as an attractor or trapping region. We refer to these settings as the local and global problems, respectively. Conditioned stochastic stability, therefore, concerns the understanding the limiting behaviour of the quasi-ergodic measure $\nu_{\varepsilon}$ for the process $X_{\varepsilon}$ conditioned upon not entering $\partial$ as $\varepsilon\to 0$.
In this paper, we show that the quasi-ergodic measure of the process $X_{\varepsilon}$ approximates, as $\varepsilon\to 0$, the equilibrium state on the maximal hyperbolic set of $T$, i.e. that of maximal topological pressure (see Section 1.3), and associated with the geometric potential $\psi=-\log|\det DT|_{E^{u}}|,$ where $E^{u}$ denotes the unstable expanding direction of $T$. This holds both locally and globally. This extends our previous results in [BassolsCastroLamb2024], where repellers are assumed to be uniformly expanding. More generally, we show that the quasi-ergodic measure for the process $X_{\varepsilon}^{\phi}$ approximates the equilibrium state associated with the potential $\psi=\phi-\log|\det DT|_{E^{u}}|$ when restricted to $\Lambda$. We dedicate the Appendix A
The rest of the manuscript is organised as follows. In Section 1.3, we briefly introduce equilibrium states, as they feature in our results when $\varepsilon\to 0$, and relate quasi-ergodic to so-called natural measures, which lie at the heart of the transient chaos literature. In Section 2 we formally introduce the systems considered (Section 2.1) and their random perturbations (Section 2.2) to state the main results of the paper in Section 2.3 (Theorems 2.6, 2.7 for the local setting and Theorems 2.8, 2.9 for the global setting). Section 3 introduces the (anisotropic) Banach spaces $B^{t,s}$ of [Baladi2018-book, Chapter 4], which we employ to establish spectral properties of the transfer operators considered. We dedicate Sections 4 and 5 to the local problem, where the random process is conditioned upon staying in a neighbourhood $V$ of a hyperbolic basic set $\Lambda$. For weights vanishing on $\partial V$, Section 4 provides the existence of the quasi-ergodic measure as well as conditioned stochastic stability in the local context. These results are extended to non-vanishing weights in Section 5. Finally, in Section 6, we prove the uniqueness of the quasi-ergodic measure supporting the hyperbolic set $\Lambda$ (Section 6.1) and promote all these results to the global setting (Section 6.2) by means of dynamical filtrations [Conley1978].
1.3. Equilibrium states and natural measures
The theory of thermodynamic formalism concerns the study of so-called equilibrium states. For a map $T:\Lambda\to\Lambda$ and a continuous weight $g:\Lambda\to\mathbb{R}$, an equilibrium state is a $T$-invariant (ergodic) Borel probability measure $\nu$ on $\Lambda$ realising the supremum of the metric pressure
$P_{\mu}(T,\Lambda,\psi)=h_{\mu}(T,\Lambda)+\int\psi\,\mathrm{d}\mu$
over all such $T$-invariant ergodic measures, where $h_{\mu}(T,\Lambda)$ denotes the Kolmogorov-Sinai (metric) entropy [Kolmogorov1958, Sinai1959] and $\psi=\log g-\log|\det DT|_{E^{u}}|$. These measures were introduced in the foundational work of Sinai, Ruelle and Bowen [Sinai1972, Bowen1975-book, BowenRuelle1975, Ruelle1976, Ruelle1978] and are motivated by ideas from the theory of statistical mechanics as they are set to minimise a theoretical analogue of the system’s free energy. Amongst equilibrium states, we distinguish those associated with the “geometric” potential $\psi=-\log|\det DT|_{E^{u}}|$, where $E^{u}$ denotes the unstable expanding direction of $T$, i.e. $g=1$. It is well-known that these equilibrium states correspond to the Sinai-Ruelle-Bowen (SRB) measure, which characterises the (natural) distribution of Lebesgue typical orbits on $\Lambda$ [Young2002].
From a different point of view, empirical results in transient chaos show that the transient dynamics of a map $T$ near a hyperbolic invariant set $\Lambda$ are governed by the statistics of nearby points in $\Lambda$ [GrebogiOttYorke1983, Lai2011]. In this context, the so-called natural measure on $\Lambda$ may be computed from the histogram of trajectories that remain close to $\Lambda$ for sufficiently long times [Kantz1985, Nusse1989, Bassols2023]. Quasi-ergodic measures, as defined above, and their limit as $\varepsilon\to 0$, aim to formalise these objects. In the process, we provide a new technique to approximate natural measures as the limit of quasi-ergodic measures, which, in turn, may be approximated from the dominant eigenfunctions of the (annealed) transfer operator studied herein.
2. Setup and results
2.1. The deterministic dynamics
Throughout this paper, unless stated otherwise, the map $T:M\to M$ is a $\mathcal{C}^{r}(M,M)$ diffeomorphism, $r\geq 1$, on an orientable compact Riemannian manifold $M$ and $g\in\mathcal{C}^{r-1}_{0}(M,\mathbb{R}^{+}_{0})$ denotes a weight function, $\mathbb{R}^{+}_{0}\coloneqq[0,\infty)$. We may assume that $M$ is embedded in an orientable boundaryless compact manifold $E$ of the same dimension and acting as an ambient space endowed with a metric compatible with that of $M$. Note that we may take $E=M$ if $M$ is itself without boundary. Denote by $\mathrm{Leb}$ a Borel measure on $E$ induced by a smooth volume form compatible with this metric, which we may think of as Lebesgue.
We consider maps $T$ for which a locally maximal hyperbolic set $\Lambda$ exists.
Definition 2.1 (Locally maximal hyperbolic set, isolating neighbourhood).
A $T$-invariant compact set $\Lambda$ is called hyperbolic if there exists an invariant decomposition $T_{\Lambda}M=E^{u}\oplus E^{s}$ of the tangent bundle over $\Lambda$ into two $DT$-invariant sub-bundles, and there exists a constant $C>0$ and $10<\rho<$ 1 such that
$\|\left.DT_{x}^{m}\right|_{E_{x}^{s}}\|\leq C\rho^{m},\qquad\|\left.DT_{x}^{-m}\right|_{E_{x}^{u}}\|\leq C\rho^{m},$
for every $m\geq 0$ and $x\in\Lambda$. The hyperbolic set $\Lambda$ is called locally maximal (or isolated) if it admits an open neighbourhood $V$ such that $\Lambda=\cap_{m\in\mathbb{Z}}T^{m}(\overline{V}).$ The set $V$ is called an isolating neighbourhood of $\Lambda$.
Definition 2.2 (Hyperbolic basic set, [Baladi2018-book, Definition 4.5]).
A hyperbolic set $\Lambda$ is called transitive if $T$ has a dense orbit in $\Lambda$. A hyperbolic basic set for $T$ is a transitive locally maximal hyperbolic set for $T$.
Notice that transitivity implies that the dimensions of $E^{u}(x)$ and $E^{s}(x)$ are constant on a hyperbolic basic set $\Lambda$. We denote these by $d_{u}$ and $d_{s}$, respectively.
Remark 2.3.
If $\Lambda$ is a locally maximal hyperbolic set but non-transitive, then it is well-known that there exists a partition of $R=\overline{\mathrm{Per}(T|_{\Lambda})}\coloneqq\overline{\{p\in\Lambda:\,p\text{ is $T$-periodic}\}}$ into finitely many non-empty compact subsets $R^{i,j},1\leq i\leq k$, $1\leq j\leq m(i)$ such that for every $i,j$:
  1. (i)
    $R^{i}=\cup_{j=1}^{m(i)}R^{i,j}$ is $T$-invariant,
  2. (ii)
    $T(R^{i,j})=R^{i,j+1\ (\mathrm{mod}\ m(i))}$,
  3. (iii)
    $T:R^{i}\to R^{i}$ is uniformly hyperbolic and topologically transitive, and
  4. (iv)
    each $T^{m(i)}:R^{i,j}\to R^{i,j}$ is uniformly hyperbolic and topologically exact.
Thus, we may always reduce to the hyperbolic basic setting when dealing with the local problem by taking a small enough isolating neighbourhood around each $R^{i}$ (see e.g. [Katok1995, Section 18.3.1]).
As explained in the Introduction, equilibrium states for Hölder weights on hyperbolic basics sets are well known to exist and to be unique [Bowen1975-book, Baladi-Tsujii2007, Gouezel-Liverani2008]. This is essential for understanding the limit of quasi-ergodic measures as $\varepsilon\to 0$, as we do not develop a thermodynamic formalism.
The so-called global problem in this paper concerns Axiom A diffeomorphisms. Let us recall some basic definitions for such maps.
Definition 2.4 (Set of non-wandering points, $NW(T)$).
A point $x\in M$ is said to be non-wandering if it admits a neighbourhood $U$ and there is $n\geq 1$ such that $T^{n}(U)\cap U\neq\emptyset$. We denote by $NW(T)$ the set of non-wandering points.
Definition 2.5 (Axiom A).
We say that a $\mathcal{C}^{r}(M)$ diffeomorphism $T:M\to M$, $r>1$, satisfies Axiom A if:
  1. (i)
    $NW(T)$ is hyperbolic, and
  2. (ii)
    the set of $T$-periodic points is dense in $NW(T)$, i.e. $NW(T)=\overline{\mathrm{Per}(T|_{\Lambda})}$.
If $T$ satisfies Axiom A, then observe that the density of periodic points in $NW(T)$ implies that $NW(T)$ is a locally maximal hyperbolic set, so $\Lambda=NW(T)$. In this setting, we further set $\Lambda_{i}\coloneqq R^{i}$ of Remark 2.3 if $T$ is not transitive.
2.2. The random perturbations
For $\varepsilon>0$, consider the random perturbations of $T$ given by $F_{\varepsilon}:[-\varepsilon,\varepsilon]^{m}\times E\to E$, where $F_{\varepsilon}(\omega,\cdot)\in\mathcal{C}^{r}(E)$ and $\partial_{\omega}F_{\varepsilon}(\omega,x)$ is surjective for all $\omega\in[-\varepsilon,\varepsilon]^{m}$. We assume that $\mathrm{dist}_{\mathcal{C}^{r}}(F_{\varepsilon}(\omega),T)\leq C\|\omega\|$ for some uniform $C>0$, where $\mathrm{dist}_{\mathcal{C}^{r}}$ denotes the $\mathcal{C}^{r}$-Whitney topology [Palis1982, Chapter 1.2]. Note that $m\geq\dim E$ by surjectivity of $\partial_{\omega}F_{\varepsilon}(\omega,x)$. Denote by $\Omega_{\varepsilon}\coloneqq([-\varepsilon,\varepsilon]^{m})^{\mathbb{Z}}$ the space of bi-infinite sequences of elements in $[-\varepsilon,\varepsilon]^{m}$ endowed with the probability measure $\mathbb{P}_{\varepsilon}\coloneqq\upsilon^{\otimes\mathbb{Z}}$, where $\upsilon$ is an absolutely continuous probability measure on $[-\varepsilon,\varepsilon]^{m}$ of full support. For $\omega\in\Omega_{\varepsilon}$, we define $T_{\omega}(x)\coloneqq T_{\omega_{0}}(x)\coloneqq F_{\varepsilon}(\omega_{0},x)$ and write $T^{n}_{\omega}(x)\coloneqq T_{\omega_{n-1}}\circ\dots\circ T_{\omega_{0}}(x)$ for every $n\in\mathbb{N}_{0}$. For $n\in\mathbb{Z},n<0$ we may write $T^{n}_{\omega}(x)=(T_{\omega}^{n})^{-1}(x).$ Moreover, we denote by $\theta$ the two-sided shift $\theta:\Omega_{\varepsilon}\to\Omega_{\varepsilon}$ with $(\theta\omega)_{i}=\omega_{i+1},i\in\mathbb{Z},$ and use $\Theta:\Omega_{\varepsilon}\times E\to\Omega_{\varepsilon}\times E$ to denote the skew-product map [^3] The symbol $\Theta$ is also used in Section 3 to denote cone systems but this will be clear by context. $(\omega,x)\mapsto(\theta\omega,T_{\omega}(x))$.
Let $X_{\varepsilon}$ be the Markov process generated by $F_{\varepsilon}$, where $\Omega=\Omega_{\varepsilon}$, $X^{\varepsilon}_{n}(\omega,x)\coloneqq T^{n}_{\omega}(x)$ for $n\in\mathbb{N}$ and $X^{\varepsilon}_{0}(\omega,x)=x$ for every $x\in M$, $\mathbb{P}_{\varepsilon}$-almost surely. Given a suitable $\phi:M\to\mathbb{R}$, we denote by $X^{\phi}_{\varepsilon}$ the $e^{\phi}$-weighted Markov process as described in the Introduction, where $\partial$ is an open subset of $M$. By abuse of notation, we say that if $T_{\omega}^{n}(x)\in\partial$ for some $n\in\mathbb{N}$, then $T_{\omega}^{n+m}(x)\in\partial$ for every $m\in\mathbb{N}.$ Observe that for such random perturbations, the annealed transfer operator associated with $X^{\phi}_{\varepsilon}$ is strong Feller (see Lemma 4.5).
2.3. Formal statement of the results
We are ready to state the main assumptions considered in this paper and the four theorems concerning the existence and the uniqueness of quasi-ergodic measures (Theorem 2.6 local, Theorem 2.8 global) and conditioned stochastic stability (Theorem 2.7, Theorem 2.9 global).
In the local setting, we shall work under the following hypothesis:
Hypothesis HL.
We say that a triple $(T,g,\Lambda)$ satisfies Hypothesis HL if:
  1. (i)
    $T:M\to M$ is a $\mathcal{C}^{r}$ diffeomorphism, $r>1$, on an open neighbourhood of $\Lambda$,
  2. (ii)
    $\Lambda$ is a basic hyperbolic set with isolating neighbourhood $V$,
  3. (iii)
    $g:M\to\mathbb{R}$ is a $\mathcal{C}^{r-1}$ weight function supported in $V$, with $g|_{\Lambda}>0$, and
  4. (iv)
    $T:\Lambda\to\Lambda$ is topologically mixing.
Assuming Hypothesis HL, there exists a unique equilibrium state $\nu^{g}$ on $\Lambda$ associated with $g$ which realises the topological pressure of the system (see [Baladi2018-book, Theorem 7.5] or [Gouezel-Liverani2008])
$P_{\mathrm{top}}(T,\Lambda,\psi)=h_{\nu^{g}}(T)+\int\psi\,\mathrm{d}\nu^{g},$
with $\psi=\log g-\log|\det DT|_{E^{u}}|$. A central object in the study of equilibrium states is the Ruelle transfer operator $\mathcal{P}_{g}$ on a suitable function space defined as
$\mathcal{P}_{g}\varphi(x)=g(x)(\varphi\circ T(x)).$
Analogously, for the randomly perturbed and $e^{\phi}$-weighted process $X_{\varepsilon}^{\phi}$, we consider the (local) annealed transfer operator on $V$ instead given by
$\mathcal{P}_{\varepsilon}\varphi(x)=e^{\phi(x)}\mathbb{E}_{\varepsilon}\left[\varphi\circ T_{\omega}(x)\cdot\mathbbm{1}_{\overline{V}}\circ T_{\omega}(x)\right].$
The following result provides the existence and uniqueness of quasi-ergodic measures around each hyperbolic basic set.
Theorem 2.6.
Assume that $(T,g,\Lambda)$ satisfies Hypothesis HL. Let $e^{\phi}=g$ on $\Lambda$. Then, for $\varepsilon>0$ small enough, $X_{\varepsilon}^{\phi}$ admits a unique quasi-ergodic measure $\nu_{\varepsilon}^{\phi}$ on $\overline{V}$ such that $\Lambda\subset\mathrm{supp}\,\nu_{\varepsilon}^{\phi}$.
Moreover, $\nu_{\varepsilon}^{\phi}(\varphi)=\mu_{\varepsilon}(\varphi\cdot g_{\varepsilon})$ for any observable $\varphi\in\mathcal{C}^{0}(\overline{V})$, with $g_{\varepsilon}\in B^{t,s}$, a suitable (anisotropic) Banach space, satisfying $\mathcal{P}_{\varepsilon}g_{\varepsilon}=r(\mathcal{P}_{\varepsilon})g_{\varepsilon}$ and $\mu_{\varepsilon}\in(B^{t,s})^{*}$ satisfying $\mathcal{P}_{\varepsilon}^{*}\mu_{\varepsilon}=r(\mathcal{P}_{\varepsilon})\mu_{\varepsilon}$.
Local conditioned stochastic stability is then given by:
Theorem 2.7 (Local conditioned stochastic stability, $\partial=M\setminus\overline{V}$).
The unique quasi-ergodic measure $\nu^{\phi}_{\varepsilon}$ on $\overline{V}$ of $X_{\varepsilon}^{\phi}$ conditioned upon not leaving $\overline{V}$ converges to $\nu^{g}$ in weak- as $\varepsilon\to 0$.
In data-driven applications and the practical analysis of transient chaos, it may not be possible to identify the locus of a hyperbolic basic set $\Lambda$ nor a suitable isolating neighbourhood $V$. Following Theorem 2.6, this information is crucial to approximate the transfer operator $\mathcal{P}_{\varepsilon}$ and compute its dominant left and right eigenfunctions, which, combined, make up the quasi-ergodic measure and approximate the underlying equilibrium state.
On the other hand, provided enough time has elapsed, a system will often settle in a trapping region $\mathcal{A}$ on which the dynamics are no longer transient. While we may not be interested in the behaviour within $\mathcal{A}$, data provides us with access to its position in the state space. It is, therefore, well-justified to ask what the statistical behaviour of the process is before reaching this global trapping region.
In Theorems 2.9 and 2.9, we address this question and show that the longest transient behaviour of the system is driven by the dynamics near the most dominant hyperbolic basic set, i.e. that of the largest topological pressure and thus of largest escape rate [DemersYoung2006]. In this context, we extend the previous Hypothesis HL to distinguish the different basic sets $\Lambda_{i},i\in\{1,\dots,k\}$ and ensure there are no cycles between them:
Hypothesis HG1.
We say that $(T,g)$ satisfies Hypothesis HG1 on $M$ if:
  1. (i)
    $T:M\to M$ is an Axiom A diffeomorphism (see Definition 2.5)
  2. (ii)
    there exists $r>1$ such that $T$ is a $\mathcal{C}^{r}(M)$ function and $g\in\mathcal{C}^{r-1}(M,\mathbb{R}^{+}_{0})$ is a non-negative weight function, [^4] This weight is global and not necessarily supported on a particular set $V$.
  3. (iii)
    $\Lambda=\cup_{i=1}^{k}\Lambda_{i}=NW(T)=\overline{\mathrm{Per}(T)}$ admits no cycles (see Definition 6.2), and
  4. (iv)
    there exists a unique basic set $\Lambda_{i}$ on which the topological pressure is maximal and realised by the unique equilibrium state $\nu^{g}$, which is supported on $\Lambda_{i}$, and
  5. (v)
    $T:\Lambda_{i}\to\Lambda_{i}$ is topologically mixing.
For this problem, we consider the (global) annealed transfer operator on $M\setminus\mathcal{A}$ for the $e^{\phi}$-weighted process $X_{\varepsilon}^{\phi}$ given by
$\widehat{\mathcal{P}}_{\varepsilon}\varphi(x)=e^{\phi(x)}\mathbb{E}_{\varepsilon}\left[\varphi\circ T_{\omega}(x)\cdot\mathbbm{1}_{M\setminus\mathcal{A}}\circ T_{\omega}(x)\right].$
The following theorem provides the existence of quasi-ergodic measures globally:
Theorem 2.8.
Assume that $(T,g)$ satisfies Hypothesis HG1. Let $e^{\phi}=g$ on $\Lambda$. Then, for $\varepsilon>0$ small enough, the following holds:
  1. (i)
    the operator $\widehat{\mathcal{P}}_{\varepsilon}$ admits a unique dominant eigenfunction $g_{\varepsilon}\in B^{t,s}$ such that $\widehat{\mathcal{P}}_{\varepsilon}g_{\varepsilon}=\lambda_{\varepsilon}g_{\varepsilon}$ with $\lambda_{\varepsilon}=r(\widehat{\mathcal{P}}_{\varepsilon})$,
  2. (ii)
    the operator $\widehat{\mathcal{P}}_{\varepsilon}^{*}$ admits a unique dominant eigenfunction $m_{\varepsilon}\in(B^{t,s})^{*}$ such that $\widehat{\mathcal{P}}_{\varepsilon}^{*}m_{\varepsilon}=\lambda_{\varepsilon}m_{\varepsilon}$ with $\lambda_{\varepsilon}=r(\widehat{\mathcal{P}}_{\varepsilon})$,
  3. (iii)
    $\nu_{\varepsilon}^{\phi}(\,\cdot\,)=m_{\varepsilon}(\cdot\,g_{\varepsilon})$ is the unique quasi-ergodic measure for $X_{\varepsilon}^{\phi}$ on $\{g_{\varepsilon}>0\}$, conditioned upon not entering $\partial=\mathcal{A}$.
Global conditioned stochastic stability is then given by:
Theorem 2.9 (Global conditioned stochastic stability, $\partial=\mathcal{A}$).
The quasi-ergodic measure $\nu^{\phi}_{\varepsilon}$ of Theorem 2.8 converges to $\nu^{g}$ of Hypothesis HG1 in weak- as $\varepsilon\to 0$.
This quasi-ergodic measure $\nu_{\varepsilon}^{\phi}$ is not necessarily unique. In the presence of finitely many basic sets $\Lambda_{1},\dots,\Lambda_{k}\subset M$, an initial condition $x$ may never feel the presence of the most dominant hyperbolic set so one should expect that $x\not\in\{g_{\varepsilon}>0\}$. Based on the classical notion of dynamical filtrations [Conley1978], and under some additional assumptions (see Hypothesis HG2), we prove a more detailed version of Theorems 2.8 and 2.9 in Section 6.2. Theorem 6.5 provides, in particular, $t\leq k$ different quasi-ergodic measures $\nu_{\varepsilon,i}^{\phi},\dots,\nu_{\varepsilon,t}^{\phi},$ and for each $x\in W^{s}(\Lambda_{j})$, for some $j\in\{1,\dots,k\}$, the quasi-ergodic limit in (1) holds for some particular $\nu_{\varepsilon,\ell}^{\phi},\ell\in\{1,\dots,t\}$. The conditioned stochastic stability of these quasi-ergodic measures is also shown.
3. The anisotropic Banach space $B^{t,s}$
This section contains a brief summary of Chapters 4 and 5 in the book “Dynamical Zeta Functions and Dynamical Determinants for Hyperbolic Maps” by Viviane Baladi [Baladi2018-book]. The construction of the anisotropic Banach spaces $B^{t,s}$ is essential to obtain bounds on the spectral and essential spectral radii of the deterministic transfer operator, which can be extended to random settings by well-established perturbation theory results [KellerLiverani1999]. Importantly, the spaces $B^{t,s}$ allow us to interpret the eigenfunctions of the dual transfer operator as (quasi-stationary) measures, which we leverage in the construction of the quasi-ergodic measure. In the process, we show a compact embedding property of these spaces (see Theorem 3.15), which is leveraged in the proof of Theorem 2.6 and Theorem 2.7 for vanishing Hölder weights (in Section 4) and in Lemma 4.4.
Notation 3.1.
We mostly follow the definitions in the underlying reference and the standard nomenclature used in the literature, where possible.
  1. (i)
    $K\subset\mathbb{R}^{d}$ is a closed subset.
  2. (ii)
    $\mathcal{C}^{\alpha}(K)$ denotes the $\alpha$-Hölder space of functions from $\mathbb{R}^{d}$ into $\mathbb{C}$, and use the subindex $\mathcal{C}^{\alpha}_{0}(K)$ to denote functions in $\mathcal{C}^{\alpha}(K)$ which vanish on the boundary of $K$, $\partial K$.
  3. (iii)
    $\mathcal{F}_{b}(Y)\coloneqq\{h:Y\to\mathbb{R};\ h\text{ is bounded and measurable}\}$ for any domain $Y\subset M$.
  4. (iv)
    $\mathcal{F}$ denotes the Fourier transform and $\mathcal{F}^{-1}$ the inverse Fourier transform.
  5. (v)
    $H^{t}_{p}(\mathbb{R}^{d})$, with $t\in\mathbb{R},1<p<\infty$ denotes the (isotropic) Sobolev space defined by
    $H^{t}_{p}(\mathbb{R}^{d})\coloneqq\{\varphi\in\mathcal{S}^{\prime}\,:\,\|\varphi\|_{H^{t}_{p}(\mathbb{R}^{d})}\coloneqq\|\mathcal{F}^{-1}\left((1+\|\xi\|^{2})^{t/2}\mathcal{F}(\varphi)(\xi)\right)\|_{L_{p}(\mathbb{R}^{d})}<\infty\},$
    where $\|\cdot\|_{L_{p}(\mathbb{R}^{d})}$ is the usual norm and $\mathcal{S}^{\prime}$ the (Schwartz) space of tempered distributions. Recall that if $\ell\in\mathbb{N}$, then
    $H^{\ell}_{p}(\mathbb{R}^{d})\equiv\{\varphi\in L_{p}(\mathbb{R}^{d})\,:\,\sum_{|\beta|\leq\ell}\|\partial^{\beta}\varphi\|_{L_{p}(\mathbb{R}^{d})}<\infty\}$
    for all $1<p<\infty$.
  6. (vi)
    Given $t>0$, we write $H^{t}$ for the $L_{2}$ Sobolev space of order $t$, i.e. $H^{t}=H^{t}_{2}$ (see [Baladi2018-book, Chapter 2, Section 2.2.1]). Recall that its dual is $H^{-\ell}$.
3.1. Charts and cone systems adapted to $(T,V)$
A cone in $\mathbb{R}^{d}$ is an invariant subset under scalar multiplication. For two cones $\mathbf{C}$ and $\mathbf{C}^{\prime}$ in $\mathbb{R}^{d}$, we write
$\mathbf{C}\Subset\mathbf{C}^{\prime}\ \text{if}\ \overline{\mathbf{C}}\subset\mathrm{Int}(\mathbf{C}^{\prime})\cup\{0\}.$
We say that a cone $\mathbf{C}$ is $d^{\prime}$-dimensional if $d^{\prime}\geq 1$ is the maximal dimension of a linear subset of $\mathbf{C}.$
Definition 3.2 (Cone system, $\Theta<\Theta^{\prime}$, [Baladi2018-book, Definition 4.10]).
Let $\mathbf{C}_{+}$ and $\mathbf{C}_{-}$ be closed cones in $\mathbb{R}^{d}$ with non-empty interiors, of respective dimensions $d_{s}$ and $d_{u}$, and such that $\mathbf{C}_{+}\cap\mathbf{C}_{-}=\{0\}$ (i.e. the cones are transversal). Let $\Phi_{+}:\mathbb{S}^{d-1}\to[0,1]$ be a $\mathcal{C}^{\infty}$ function on the unit sphere $\mathbb{S}^{d-1}$ in $\mathbb{R}^{d}$ satisfying
$\Phi_{+}(\xi)=\begin{cases}1,&\xi\in\mathbb{S}^{d-1}\cap\mathbf{C}_{+}\\ 0,&\xi\in\mathbb{S}^{d-1}\cap\mathbf{C}_{-}.\end{cases}$
Define $\Phi_{-}:\mathbb{S}^{d-1}\to[0,1]$ by $\Phi_{-}(\xi)=1-\Phi_{+}(\xi).$ A quadruple $\Theta=(\mathbf{C}_{+},\mathbf{C}_{-},\Phi_{+},\Phi_{-})$ is called a cone system. For another such quadruple $\Theta^{\prime}=(\mathbf{C}_{+}^{\prime},\mathbf{C}_{-}^{\prime},\Phi_{+}^{\prime},\Phi_{-}^{\prime}),$ we write $\Theta<\Theta^{\prime}$ if $\mathbb{R}^{d}\setminus\mathbf{C}^{\prime}_{+}\Subset\mathbf{C}_{-}.$ In particular, this implies $\mathbf{C}_{+}\Subset\mathbf{C}_{+}^{\prime}$ and $\mathbf{C}_{-}^{\prime}\Subset\mathbf{C}_{-}$.
Definition 3.3 (Cone-hyperbolic diffeomorphism, [Baladi2018-book, Definition 4.11]).
Let $U$ be an open and bounded set in $\mathbb{R}^{d}$, let $\Theta=(\mathbf{C}_{+},\mathbf{C}_{-},\Phi_{+},\Phi_{-})$ and $\Theta^{\prime}=(\mathbf{C}_{+}^{\prime},\mathbf{C}_{-}^{\prime},\Phi_{+}^{\prime},\Phi_{-}^{\prime})$ be two cone systems. A $\mathcal{C}^{r}$ diffeomorphism $F:U\to\mathbb{R}^{d}$ onto its image is cone-hyperbolic from $\Theta$ to $\Theta^{\prime}$ if $F$ extends to a bilipschitz $\mathcal{C}^{1}$ diffeomorphism of $\mathbb{R}^{d}$ such that
$DF^{\intercal}_{x}(\mathbb{R}^{d}\setminus\mathbf{C}_{+})\Subset\mathbf{C}^{\prime}_{-}\ \text{for every }x\in\mathbb{R}^{d},$
where $A^{\intercal}$ denotes the transpose of the matrix $A$.
Definition 3.4 (Charts and partition of unity adapted to $(T,V)$, [Baladi2018-book, Definition 4.14]).
Fix a finite system of $\mathcal{C}^{\infty}$ local charts $\{(V_{i},\kappa_{i})\}_{i\in I},$ with open subsets $V_{i}\subset M$ and maps $\kappa_{i}:V_{i}\to\mathbb{R}^{d}$ such that $V\subset\cup_{i}V_{i}$ and
  1. (1)
    $\mathcal{V}=\{V_{i}\}_{i\in I}$ is a generating cover [^5] Recall that a cover is generating for $T$ if the diameter of all sets $\{\cap_{k=0}^{m-1}T^{-k}(V_{\omega_{k}}):\,\omega\in I^{m}\}$ tends to zero as $m\to\infty$. of $V$, and there is no strict sub-cover.
  2. (2)
    $\kappa_{i}(V_{i})$ is a bounded open subset of $\mathbb{R}^{d}$ with smooth boundary for each $i\in I$.
Finally, let $\{\theta_{i}\}_{i\in I}$ be a $\mathcal{C}^{\infty}$ finite partition of unity for $V$ subordinate to the cover $\mathcal{V}$, that is, the support of each $\theta_{i}:M\to[0,1]$ is contained in $V_{i}$, and we have $\sum_{i\in I}\theta_{i}(x)=1$ for every $x\in V.$
Definition 3.5 (Cone systems adapted to $(T,V)$, [Baladi2018-book, Definition 4.15]).
If $\Lambda$ is a hyperbolic basic set for $T$ with isolating neighbourhood $V$, we may choose a finite family of cone systems $\{\Theta_{i}=(\mathbf{C}_{i,+},\mathbf{C}_{i,-},\Phi_{i,+},\Phi_{i,-}\}_{i\in I}$, where $I$ is the index set for the local charts from Definition 3.4, so that the following conditions hold:
  1. (3)
    If $x\in V_{i}\cap\Lambda$, the cone $(D\kappa_{i})^{*}_{x}(\mathbf{C}_{i,+})$ contains the ($d_{s}$-dimensional) normal subspace of $E^{u}(x)$, and the cone $(D\kappa_{i})^{*}_{x}(\mathbf{C}_{i,-})$ contains the ($d_{u}$-dimensional) normal subspace of $E^{s}(x).$
  2. (4)
    If $V_{ij}\coloneqq T^{-1}(V_{j})\cap V_{i}\neq\emptyset$, the map in charts
    $F=T_{ij}=\kappa_{j}\circ T\circ\kappa_{i}^{-1}:\kappa_{i}(V_{ij})\to\kappa_{j}(V_{j})$
    is a $\mathcal{C}^{r}$ cone-hyperbolic diffeomorphism from $\Theta_{i}$ to $\Theta_{j}$.
3.2. The space $B^{\Theta,t,s}$ in $\mathbb{R}^{d}$
We follow the presentation given in [Baladi2018-book, Section 2.4.1]. Let $\chi:\mathbb{R}\to[0,1]$ be a symmetric $\mathcal{C}^{\infty}$ bump-function such that
$\chi(x)=\begin{cases}1,&|x|\leq 1,\\ 0,&|x|\geq 2.\end{cases}$
Define the functions $\psi_{n}:\mathbb{R}^{d}\to\mathbb{R},n\in\mathbb{N}$, such that $\psi_{0}(\xi)=\chi(\|\xi\|)$ and
$\psi_{n}(\xi)=\chi(2^{-n}\|\xi\|)-\chi(2^{-n+1}\|\xi\|),\text{ for }n\geq 1.$
Given a cone system $\Theta=(\mathbf{C}_{+},\mathbf{C}_{-},\Phi_{+},\Phi_{-})$ in $\mathbb{R}^{d}$, let $(n,\sigma)\in(\mathbb{N}\cup\{0\})\times\{-,+\}$. We define for $\xi\in\mathbb{R}^{d}:$
$\psi_{\Theta,0,\sigma}(\xi)\coloneqq\frac{\chi(\|\xi\|)}{2},\quad\text{ and }\quad\psi_{\Theta,n,\sigma}(\xi)\coloneqq\Phi_{\sigma}\left(\frac{\xi}{\|\xi\|}\right)\psi_{n}(\xi),\text{ for }\ n\geq 1.$
Observe that
$\sum_{\sigma\in\{+,-\}}\sum_{n=0}^{\infty}\psi_{\Theta,n,\sigma}=1.$
Notation 3.6.
Given a Hölder function $\varphi$ we define $\varphi_{\Theta,n,\sigma}$ to be
$\varphi_{\Theta,n,\sigma}\coloneqq\widehat{\psi}_{\Theta,n,\sigma}\ast\varphi\coloneqq\int_{\mathbb{R}^{d}}\widehat{\psi}_{\Theta,n,\sigma}(x-y)\varphi(y)\mathrm{d}y,$
where
$\widehat{\psi}_{\Theta,n,\sigma}(x)\coloneqq\mathcal{F}^{-1}{\psi}_{\Theta,n,\sigma}(x)\coloneqq\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}e^{ix\xi}\psi_{\Theta,n,\sigma}(\xi)\mathrm{d}\xi,$
with $x\xi$ denoting the scalar product.
Definition 3.7 (Fake unstable leaves $\mathcal{F}(\Theta)$ and the $L_{1}(\mathcal{F})$-norm, [Baladi2018-book, Definition 5.10]).
Let $\Theta=(\mathbf{C}_{+},\mathbf{C}_{-},\Phi_{+},\Phi_{-})$ be a cone system. We define the set of unstable leaves $\mathcal{F}=\mathcal{F}(\Theta)$ to be the set of $\mathcal{C}^{1}$-submanifolds $\Gamma\subset\mathbb{R}^{d}$, of dimension $d_{u}$, such that the straight line connecting any two points in $\Gamma$ is normal to a $d_{s}$-dimensional subspace contained in $\mathbf{C}_{+}$, i.e.
$\sup_{\begin{subarray}{c}x,y\in\Gamma\\ x\neq y\end{subarray}}\inf_{\begin{subarray}{c}C\subset\mathbf{C}_{+}\\ d_{s}-\text{subspace}\end{subarray}}\sup_{c\in C\setminus\{0\}}\left\langle\frac{y-x}{\|y-x\|},\frac{c}{\|c\|}\right\rangle=0.$
For $\varphi\in\mathcal{C}^{\infty}(\mathbb{R}^{d})$ and $\mathcal{F}=\mathcal{F}(\Theta)$ we introduce the $L_{1}(\mathcal{F})$-norm
$\|\varphi\|_{L_{1}(\mathcal{F})}=\sup_{\Gamma\in\mathcal{F}}\|\varphi\|_{L_{1}(\mu_{\Gamma})},$
where $\mu_{\Gamma}$ is the Riemannian volume on $\Gamma$ induced by the standard metric on $\mathbb{R}^{d}$.
Definition 3.8 (The local space $B^{\Theta,t,s}(K)$, [Baladi2018-book, Definition 5.12]).
Given real numbers $t,s$, a cone system $\Theta$, and a compact set $K\subset\mathbb{R}^{d}$, we define
$\|\varphi\|_{B^{\Theta,t,s}}\coloneqq\max\left\{\sup_{n\geq 0}2^{tn}\|\varphi_{\Theta,n,+}\|_{L_{1}(\mathcal{F})},\sup_{n\geq 0}2^{sn}\|\varphi_{\Theta,n,-}\|_{L_{1}(\mathcal{F})}\right\}$
and set $B^{\Theta,t,s}$ to be the completion of $\mathcal{C}^{u}_{0}(K)$, for any fixed $u>t$, with respect to the $B^{\Theta,t,s}$-norm, $\|\cdot\|_{B^{\Theta,t,s}}$.
Proposition 3.9 ([Baladi2018-book, Lemma 5.14]).
Assume $s\leq t$. For any $u>\max\{0,t\}$, there exists a constant $C=C(u,K)$ such that $\|\varphi\|_{B^{\Theta,t,s}}\leq C\|\varphi\|_{\mathcal{C}^{u}_{0}}$ for every $\varphi\in\mathcal{C}^{\infty}(\mathbb{R}^{d};\mathbb{C})$ supported in $K$. Moreover, for any $v>\max\{-s,0\}$, the space $B^{\Theta,t,s}(K)$ is contained in the space of distributions of order $v$ supported in $K$.
In other words, $\mathcal{C}^{u}_{0}(K)\subset B^{\Theta,t,s}\subset(\mathcal{C}^{v}_{0}(K))^{*}$ if $u>\max\{0,t\}$ and $v>\max\{-s,0\}$.
Notation 3.10.
Consider $K\subset\mathbb{R}^{d}$ a compact subset with smooth boundary, a finite open cover $\mathcal{V}=\{V_{i}\}_{i\in I}$, a system of $\mathcal{C}^{\infty}$ local charts $\{\kappa_{i}:V_{i}\to\mathbb{R}^{d}\}_{i\in I}$ and a $\mathcal{C}^{\infty}$ partition of unity $\{\theta_{i}:K\to[0,1]\}_{i\in I}$ subordinate to $\mathcal{V}$. For $t\geq 0$ and $1<p<\infty$, given $\varphi:M\to\mathbb{C}$, we define the $H^{t}_{p}(K)$-norm to be
$\|\varphi\|_{H^{t}_{p}(K)}\coloneqq\max_{i\in I}\|(\theta_{i}\cdot\varphi)\circ\kappa_{i}^{-1}\|_{H^{t}_{p}(\mathbb{R}^{d})},$
and set $H^{t}_{p}(K)\coloneqq\{\varphi\in L_{p}(K)\,:\,\|\varphi\|_{H^{t}_{p}(K)}<\infty\}.$
Corollary 3.11.
Assume $s\leq t$. Suppose that $K\subset\mathbb{R}^{d}$ is a compact set with smooth boundary, then there exists $\ell>0$ such that the Banach space $B^{\Theta,t,s}(K)$ compactly embeds into $H^{-\ell}(K)$.
Proof.
It follows from a general Sobolev inequality (see e.g. [Evans2010, Section 5.6.3, Theorem 6]) that $H^{\ell}(K)$, for $\ell>d/2$ an integer, compactly embeds into $\mathcal{C}^{v}_{0}(K)$ for any $v<\ell-(d/2)_{+}$, where
$x_{+}\coloneqq\begin{cases}x,&x\not\in\mathbb{N}\\ x+\delta,&x\in\mathbb{N},\end{cases}$
with $10<\delta<$ 1 arbitrary. Taking the dual of this embedding we obtain that $(\mathcal{C}^{v}_{0}(K))^{*}$ compactly embeds into $(H^{\ell}(K))^{*}=H^{-\ell}(K)$. We may choose $\ell>d/2$ large enough such that $\ell>\ell-(d/2)_{+}>v>\max\{-s,0\}$ and conclude using Proposition 3.9. ∎
3.3. The space $B^{t,s}$ on $M$
The ideas developed so far can be adapted to $M$ and the map $T$ by means of suitable local charts.
Definition 3.12 (Regular cone-hyperbolicity, [Baladi2018-book, Definition 5.10]).
Let $U$ be a bounded open subset in $\mathbb{R}^{d}$, and let $\Theta=(\mathbf{C}_{+},\mathbf{C}_{-},\Phi_{+},\Phi_{-})$ and $\Theta^{\prime}=(\mathbf{C}_{+}^{\prime},\mathbf{C}_{-}^{\prime},\Phi_{+}^{\prime},\Phi_{-}^{\prime})$ be two cone systems. A $\mathcal{C}^{r}$ diffeomorphism onto its image $F:U\to\mathbb{R}^{d}$ is regular cone-hyperbolic from $\Theta$ to $\Theta^{\prime}$ if $F$ is cone-hyperbolic in the sense of Definition 3.3 and, in addition, there exists, for each $x,y\in U$, a linear transformation $\mathbb{L}_{xy}$ satisfying
$\mathbb{L}_{x,y}^{\intercal}(\mathbb{R}^{d}\setminus\mathbf{C}_{+})\Subset\mathbf{C}_{-}^{\prime}\text{ and }\mathbb{L}_{xy}(x-y)=F(x)-F(y).$
Notice that if $F$ is regular cone-hyperbolic, then the extension of $F$ to $\mathbb{R}^{d}$ maps each element of $\mathcal{F}(\Theta^{\prime})$ to an element of $\mathcal{F}(\Theta)$.
Definition 3.13 (Anisotropic spaces $B^{t,s}$ on $M$).
Fix real numbers $s$ and $t$. The Banach space $B^{t,s}=B^{\Theta,t,s}(T,V)$ is the completion of $\mathcal{C}^{\infty}(\overline{V})$ for the norm
$\|\varphi\|_{B^{t,s}(T,V)}\coloneqq\max_{i\in I}\|(\theta_{i}\cdot\varphi)\circ\kappa_{i}^{-1}\|_{B^{\Theta_{i},t,s}},$
where $\{\kappa_{i}:V_{i}\to\mathbb{R}^{d}\}_{i\in I}$ is a finite $\mathcal{C}^{\infty}$ atlas of $M$, $\{\theta_{i}\}_{i\in I}$ is a partition of unity and $\Theta=\{\Theta_{i}\}_{i\in I}$ a family of cone systems, which satisfy the requirements of Definitions 3.4 and 3.5, where item 4 in Definition 3.5 is strengthened to require regular cone-hyperbolicity as in Definition 3.12.
Observe that the symbol $\Theta$ refers to a single cone system when considering $B^{\Theta,t,s}(K)$ for some compact $K\subset\mathbb{R}^{d}$ but refers to a finite family of cone systems, satisfying the regular cone-hyperbolic conditions of Definitions 3.3 and 3.12 when considering $B^{\Theta,t,s}(T,V)$. Moreover, we may drop the symbol $\Theta$ and simply write $B^{t,s}$ when a statement holds for any family of cone systems.
Proposition 3.14 (Leibniz bounds on $B^{t,s}(T,V)$).
Consider $r>0$ and $t-(r-1)<s<t$, a generating open cover $\mathcal{V}=\{V_{i}\}_{i\in I}$ of $M$, a finite smooth atlas $\{\kappa_{i}:V_{i}\to\mathbb{R}^{d}\}_{i\in I}$ of $M$, a partition of unity $\{\theta_{i}\}_{i\in I}$, and two finite families of cone systems in $\mathbb{R}^{d}$, $\Theta=\{\Theta_{i}\}_{i\in I}$ and $\Theta^{\prime}=\{\Theta_{i}^{\prime}\}_{i\in I}$, such that $\Theta_{i}^{\prime}<\Theta_{i}$ for every $i\in I$, all satisfying the conditions of Definition 3.4. Then, there exists $C=C(\Theta,\Theta^{\prime})$ such that
$\|\varphi\|_{B^{\Theta^{\prime},t,s}(T,V)}\leq C\|\varphi\|_{B^{\Theta,t,s}(T,V)}$
for every $\varphi\in B^{\Theta,t,s}(T,V).$
Proof.
For each $i\in I$, choose a $\mathcal{C}^{\infty}$ function function $\widetilde{\theta}_{i}:M\to[0,1]$ such that
$\widetilde{\theta}_{i}(x)=\begin{cases}1,&x\in\mathrm{supp}(\theta_{i})\Subset V_{i},\\ 0,&x\in M\setminus V_{i}.\end{cases}.$
From [Baladi2018-book, Corollary 5.19] we obtain that for every $i\in I$, there exists $C_{i}=C_{i}(\Theta,\Theta^{\prime})$ such that
$\|\left(\theta_{i}\cdot\varphi\right)\circ\kappa_{i}^{-1}\|_{B^{\Theta_{i}^{\prime},t,s}(\overline{V_{i}})}$ $=\|(\widetilde{\theta}_{i}\circ\kappa_{i}^{-1})\cdot\left(\theta_{i}\cdot\varphi\right)\circ\kappa_{i}^{-1}\|_{B^{\Theta_{i},t,s}}$
$\leq C_{i}\|\widetilde{\theta}_{i}\circ\kappa_{i}^{-1}\|_{\mathcal{C}^{r-1}}\|\left(\theta_{i}\cdot\varphi\right)\circ\kappa_{i}^{-1}\|_{B^{\Theta_{i}^{\prime},t,s}(\overline{V_{i}})}.$
Set $C=\max_{i\in I}C_{i}\|\widetilde{\theta}_{i}\circ k_{i}^{-1}\|_{\mathcal{C}^{r-1}}$. ∎
Below, we state the main result of this section. Although the following theorem appears natural, it does not appear to have been previously established in the literature (see [Baladi-Tsujii2007, Baladi-Tsujii2008, Baladi2018-book]). [^6] This is likely because quasi-compactness, in the cited works, is not derived via a standard Lasota-Yorke-type inequality. Instead, the approach relies on decomposing the transfer operator $\mathcal{L}_{g}$ as $\mathcal{L}_{g}=\mathcal{L}_{c}+\mathcal{L}_{b},$ where $\mathcal{L}_{c}:\mathcal{B}^{t,s}\to\mathcal{B}^{t,s}$ is a compact operator, and $\mathcal{L}_{b}:\mathcal{B}^{t,s}\to\mathcal{B}^{t,s}$ is a bounded operator with spectral radius strictly smaller than that of $\mathcal{L}_{g}$, i.e. $r(\mathcal{L}_{b})<r(\mathcal{L}_{g})$.
Theorem 3.15 (Compact embedding of $B^{t,s}$).
Let $s,s^{\prime}<0$, $t,t^{\prime}>0$ and consider $\Theta^{\prime}=\{\Theta_{i}^{\prime}\}_{i\in I},\Theta=\{\Theta_{i}\}_{i\in I}$ two finite families of cone systems in $\mathbb{R}^{d}$. Suppose that $t^{\prime}>t$, $s^{\prime}>s$ and $\Theta_{i}<\Theta_{i}^{\prime}$ for all $i\in I$. Assume that $s\leq t$ and $s^{\prime}\leq t^{\prime}$ then $B^{\Theta^{\prime},t^{\prime},s^{\prime}}$ compactly embeds in $B^{\Theta,t,s}.$
Proof.
Observe that it is enough to show that $B^{\Theta^{\prime},t^{\prime},s^{\prime}}(K)$ compactly embeds into $B^{\Theta,t,s}(K)$ for a compact domain $K\subset\mathbb{R}^{d}$ with smooth boundary. Given $\varepsilon>0$, take $N=N(\varepsilon)$ such that $1/2^{n\min\{t^{\prime}-t,s^{\prime}-s\}}<\varepsilon$ for all $n\geq N$. From Proposition 3.14 and the definition of the ${B^{\Theta,t,s}}$-norm we obtain that for every $\varphi\in B^{\Theta^{\prime},t^{\prime},s^{\prime}}(K)$,
$\|\varphi\|_{B^{\Theta,t,s}(K)}$ $\leq\max_{\begin{subarray}{c}n\in\{0,\ldots,N-1\}\end{subarray}}\left\{2^{tn}\|\varphi_{\Theta,n,+}\|_{L_{1}(\mathcal{F})},2^{sn}\|\varphi_{\Theta,n,-}\|_{L_{1}(\mathcal{F})}\right\}$
$\qquad+\frac{1}{2^{N\min\{t^{\prime}-t,s^{\prime}-s\}}}\|\varphi\|_{B^{\Theta,t^{\prime},s^{\prime}}}$
$\leq C_{\varepsilon}\max_{\begin{subarray}{c}n\in\{0,\ldots,N\}\\ \sigma\in\{-,+\}\end{subarray}}\{\|\varphi_{\Theta,n,\sigma}\|_{L_{1}(\mathcal{F})}\}+\varepsilon C\|\varphi\|_{B^{\Theta^{\prime},t^{\prime},s^{\prime}}(K)}.$
for some constants $C,C_{\varepsilon}>0$. To prove the theorem, it is enough to show that for every $n\in\mathbb{N}\cup\{0\}$ and $\sigma\in\{-,+\}$ the map
$F_{n,\sigma}:B^{\Theta,t,s}(K)$ $\to L_{1}(\mathcal{F})$
$\varphi$ $\mapsto\widehat{\psi}_{\Theta,n,\sigma}\ast\varphi.$
is a compact linear operator. In fact, from Corollary 3.11, we may choose $\ell\in\mathbb{N}$ large enough such that the inclusion $\iota:B^{\Theta,t,s}\to H^{-\ell}(K)$ is a compact linear operator. Thus, it suffices to show that
$G_{n,\sigma}:H^{-\ell}(K)$ $\to L_{1}(\mathcal{F})$
$\varphi$ $\mapsto\widehat{\psi}_{\Theta,n,\sigma}\ast\varphi$
is a bounded linear operator. Recall that the function $\widehat{\psi}_{\Theta,n,\sigma}$ lies in the Schwartz class for every $(n,\sigma)\in\mathbb{N}_{0}\times\{-,+\}.$ It follows that given $\varphi\in\mathcal{C}^{\infty}(K)$ and $k\in\mathbb{N}$ such that $\int_{\mathbb{R}^{d}}1/(1+\|x\|^{k})\mathrm{d}x<\infty$, for any compactly supported smooth function $\overline{\theta}:\mathbb{R}^{d}\to[0,1]$ such that $\bar{\theta}(x)=1$ for every $x\in K$, we obtain that
$|\widehat{\psi}_{\Theta,n,\sigma}\ast\varphi(x)|$ $=\left|\int_{\mathbb{R}^{d}}\varphi(y)\widehat{\psi}_{\Theta,n,\sigma}(x-y)\mathrm{d}y\right|$
$\leq\left|\int_{K}\varphi(y)\overline{\theta}(y)\widehat{\psi}_{\Theta,n,\sigma}(x-y)\mathrm{d}y\right|$
$\leq\|\varphi\|_{H^{-\ell}(K)}\|\overline{\theta}(\,\cdot\,)\widehat{\psi}_{\Theta,n,\sigma}(x-\cdot)\|_{H^{\ell}(\mathbb{R}^{d})}.$ (4)
Recall that for $\ell\in\mathbb{N}$, we have that
$\|\overline{\theta}(\cdot)\widehat{\psi}_{\Theta,n,\sigma}(x-\cdot)\|^{2}_{H^{\ell}(\mathbb{R}^{d})}=\sum_{|\beta|\leq\ell}\int_{\mathbb{R}^{d}}\left|D^{\beta}\left(\overline{\theta}(y)\widehat{\psi}_{\Theta,n,\sigma}(x-y)\right)\right|^{2}\mathrm{d}y.$ (5)
The derivative operator sends compactly supported functions to compactly supported functions and likewise for functions in the Schwartz class $\mathcal{S}$. Moreover, for each $\psi\in S$ and for each $k\in\mathbb{N}$, there exists $C_{k}>0$ such that $|\psi(x)|^{2}\leq C_{k}/(1+\|x\|^{k})$. This provides an integrable bound to the right-hand side of (5) for every multi-index $\beta$, i.e. combining this argument with (3.3) and (several) triangle inequalities we obtain
$\begin{split}\|\widehat{\psi}_{\Theta,n,\sigma}\ast\varphi\|_{L_{1}(\mathcal{F})}&\leq\|\varphi\|_{H^{-\ell}(K)}\cdot\|\|\overline{\theta}(\,\cdot\,)\widehat{\psi}_{\Theta,n,\sigma}(x-\cdot)\|_{H^{\ell}(\mathbb{R}^{d})}\|_{L_{1}(\mathcal{F})}\\ &\leq\|\varphi\|_{H^{-\ell}(K)}\sup_{\Gamma\in L_{1}(\mathcal{F})}\int_{\Gamma}\frac{\widetilde{C}_{k}}{1-\|x\|^{k}}\mathrm{d}\mu_{\Gamma}(x)\leq K\|\varphi\|_{H^{-\ell}(K)},\end{split}$
under a suitable choice of $k$ large and for some constants $\widetilde{C}_{k},K>0$, where we have used the bound
$\int_{\mathbb{R}^{d}}|\bar{\theta}(y)\psi(x-y)|^{2}\mathrm{d}y\leq\int_{\mathrm{supp}\,\overline{\theta}}\bar{\theta}^{2}(y)\frac{C_{k_{0}}^{2}}{1+\|x-y\|^{2k_{0}}}\mathrm{d}y\leq\widetilde{C}_{k_{0}}\frac{1}{1+\|x\|^{2k_{0}}}$
to conclude. ∎
4. Properties of $\mathcal{P}_{g}$ on $B^{t,s}$ and local results for vanishing weights
We dedicate this section to the study of the deterministic transfer operator $\mathcal{P}_{g}$ and the annealed operator $\mathcal{P}_{\varepsilon}$ acting on $B^{t,s}$ for Hölder weights $\psi:V\to\mathbb{R}$ vanishing on $\partial V$. In particular, we present a strategy to build quasi-ergodic measures in Lemma 4.2 based on the presence of a spectral gap in $\mathcal{C}^{0}$ and check that the required assumptions are satisfied.
Let us first recall some standard definitions from operator theory.
Definition 4.1 (Quasi-compactness, spectral gap).
Let $B$ be a Banach space and $L:B\to B$ be a bounded linear transformation. We say that $L$ is quasi-compact if there exist $L$-invariant sub-spaces $F,W\subset B$ such that $F\otimes W=B$, $F$ is finite-dimensional and $r(\left.T\right|_{W})<r(T)$. We say that $L$ has a spectral gap if $L$ is a quasi-compact operator and there exists only one element in the peripheral spectrum of $L$, and such element is a simple eigenvalue.
Lemma 4.2.
Let $(\Omega,\mathbb{P})$ be a probability space and $\{T_{\omega}:M\to M\}_{\omega\in\Omega}$ be a family of maps. For $Y\subset M$ a compact subset and a non-negative function $\psi:Y\to\mathbb{R}$, $\psi\in\mathcal{C}^{\alpha}(Y)$, $\alpha>0$, assume that the operator
$\mathcal{P}:f(x)\mapsto\psi(x)\mathbb{E}_{\varepsilon}[f\circ T_{\omega}(x)\cdot\mathbbm{1}_{Y}\circ T_{\omega}(x)]$
is strong Feller and has a spectral gap in $\mathcal{C}^{0}(Y)$. Let $g\in\mathcal{C}^{0}(Y)$ be the dominant eigenfunction such that $\mathcal{P}g=\lambda g$ with $\lambda=r(\mathcal{P})$, and let $\mu\in\mathcal{M}(Y)$ satisfy $\mathcal{P}^{*}\mu=\lambda\mu$. Assume that $\mu(g)=1$. Then, the following holds:
  1. (1)
    The measure $\nu\in\mathcal{M}(Y)$, defined by $\nu(h)\coloneqq\mu(h\cdot g)$ for every $h\in\mathcal{F}_{b}(Y)$, bounded and measurable, is a quasi-ergodic measure for the process $X^{\log\psi}$ conditioned upon staying in $Y$.
  2. (2)
    For every $x\in\{g\neq 0\}$ and every $h:Y\to\mathbb{R}$ bounded and measurable we have that
    $\lim_{n\to\infty}\mathbb{E}_{x}^{\log\psi}\left[\frac{1}{n}\sum_{i=0}^{n-1}h\circ X_{i}^{\log\psi}\,\bigg{|}\,\tau>n\right]=\int_{Y}h\mathrm{d}\nu.$
  3. (3)
    Assume that $(\Omega,\mathbb{P})=(\Omega_{\varepsilon},\mathbb{P}_{\varepsilon})$ and let $T_{\omega}$ bed defined as in Section 2.2. Then, there exists $m\in\mathcal{C}^{0}(Y)$ such that $\mu=m(x)\mathrm{d}x$ and
    $m(x)=\frac{1}{\lambda}\mathbb{E}_{\varepsilon}\left[\mathbbm{1}_{Y}\circ T_{\omega}^{-1}\frac{\psi\circ T_{\omega}^{-1}(x)}{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}m\circ T_{\omega}^{-1}(x)\right].$
Proof.
The proofs of items (1) and (2) are based on [BassolsCastroLamb2024, Appendix A] and are included here for the sake of completeness. We first prove (2), which in turn implies implies (1). Let $x\in\{g\neq 0\}$ and $h:Y\to\mathbb{R}$ be a bounded and measurable observable. We compute the conditioned Birkhoff average as follows:
$\mathbb{E}_{x}^{\log\psi}\left[\frac{1}{n}\sum_{i=0}^{n-1}h\circ X_{i}^{\log\psi}\,\bigg{|}\,\tau>n\right]$ $=\frac{1}{\mathcal{P}^{n}(\mathbbm{1}_{Y})}\frac{1}{n}\sum_{i=0}^{n-1}\mathcal{P}^{i}(h\mathcal{P}^{n-i}(\mathbbm{1}_{Y}))$
$=\frac{\lambda^{n}}{\mathcal{P}(\mathbbm{1}_{Y})}\frac{1}{n}\sum_{i=0}^{n-1}\frac{1}{\lambda^{i}}\mathcal{P}^{i}\left(h\frac{1}{\lambda^{n-i}}\mathcal{P}^{n-i}(\mathbbm{1}_{Y})\right)$
$=\frac{\lambda^{n}}{\mathcal{P}^{n}(\mathbbm{1}_{Y})}\frac{1}{n}\left[\sum_{i=0}^{n-1}\frac{1}{\lambda^{i}}\mathcal{P}^{i}\left(h\left(\frac{1}{\lambda^{n-i}}\mathcal{P}^{n-i}\mathbbm{1}_{Y}-g\cdot\mu(Y)\right)\right)\right.$
$\left.\hskip 113.81102pt+\mu(Y)\sum_{i=0}^{n-1}\frac{1}{\lambda^{i}}\mathcal{P}^{i}\left(hg\right)\right],$
where the last step consists of adding and subtracting $g\cdot\mu(Y)$, bringing the sum inside and rearranging the indexes. Since
  • $\lim_{n\to\infty}\left\|\frac{1}{\lambda^{n}}\mathcal{P}^{n}\mathbbm{1}_{Y}-g\cdot\mu(Y)\right\|_{\infty}=0$, converging exponentially fast,
  • there exists $C>0$ such that $\left\|\frac{1}{\lambda^{k}}\mathcal{P}^{k}f\right\|_{\infty}\leq C$ for every $k\in\mathbb{N}$ and every bounded and measurable function $f:M\to\mathbb{R}$, and
  • the operator $\mathcal{P}$ has a spectral gap and hence it is a mean ergodic operator satisfying [Eisner2015-book, Castro2023]
    $\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\frac{1}{\lambda^{i}}\mathcal{P}^{i}\left(hg\right)(x)=g(x)\int hg\mathrm{d}\mu,$
we obtain that
$\lim_{n\to\infty}\mathbb{E}_{x}^{\log\psi}\left[\frac{1}{n}\sum_{i=0}^{n-1}h\circ X_{i}^{\log\psi}\,\bigg{|}\,\tau>n\right]$ $=\frac{1}{g(x)\mu(Y)}\left(0+g(x)\mu(Y)\int_{Y}hg\mathrm{d}\mu\right)$
$=\frac{g(x)}{g(x)}\frac{\mu(Y)}{\mu(Y)}\nu(h)=\nu(h).$
It follows that $\nu$ is a quasi-ergodic measure and the conditioned Birkhoff averages converge for all $x\in\{g\neq 0\}$.
Finally, we show (3). From the choice of the random perturbation introduced in Section 2.2, we obtain that $\mathcal{P}^{*}\delta_{x}(\mathrm{d}y)\ll\mathrm{Leb}_{Y}(\mathrm{d}y)$ for every $x\in Y$, where $\delta_{x}(\mathrm{d}y)$ is the Dirac measure at $x$. Moreover, since
$\int_{Y}\mathcal{P}^{*}\delta_{x}(\mathrm{d}y)\mu(\mathrm{d}x)=\lambda\mu(\mathrm{d}y),$
we obtain that ${\mu(\mathrm{d}y)\ll\mathrm{Leb}_{Y}(\mathrm{d}y).}$ In this way, there exists a density $m\in L_{1}(Y,\mathrm{Leb}_{Y})$ such that $\mu(\mathrm{d}x)=m(x)\mathrm{d}x.$ Observe that for each $f\in\mathcal{C}^{0}(Y)$ we obtain that
$\lambda\int_{Y}f\mathrm{d}\mu$ $=\int_{Y}\mathcal{P}f(x)m(x)\mathrm{d}x=\int_{Y}\mathbb{\psi}(x)\mathbb{E}_{\varepsilon}\left[f\circ T_{\omega}(x)\cdot\mathbbm{1}_{Y}\circ T_{\omega}(x)\right]m(x)\mathrm{d}x$
$=\mathbb{E}_{\varepsilon}\left[\int_{T_{\omega}(Y)}f(x)\mathbbm{1}_{Y}(x)\psi\circ T_{\omega}^{-1}(x)\frac{m\circ T_{\omega}^{-1}(x)}{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}\mathrm{d}x\right]$
$=\int_{Y}f(x)\mathbb{E}_{\varepsilon}\left[\frac{\psi\circ T_{\omega}^{-1}(x)}{{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}}m\circ T_{\omega}^{-1}(x)\cdot\mathbbm{1}_{Y}\circ T_{\omega}^{-1}(x)\right]\mathrm{d}x.$
The above equation implies that
$m(x)=\frac{1}{\lambda}\mathbb{E}_{\varepsilon}\left[\frac{\psi\circ T_{\omega}^{-1}(x)}{{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}}m\circ T_{\omega}^{-1}(x)\cdot\mathbbm{1}_{Y}\circ T_{\omega}^{-1}(x)\right].$ (6)
We check that $m_{\varepsilon}\in\mathcal{C}^{0}(Y)$. From (6) we obtain that that for every $x\in Y$
$m(x)$ $\leq\sup_{(\omega,x)\in\Omega\times Y}\left|\frac{\psi\circ T_{\omega}^{-1}(x)}{{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}}\right|\mathbb{E}_{\varepsilon}[m\circ T_{\omega}^{-1}(x)\cdot\mathbbm{1}_{Y}\circ T_{\omega}^{-1}(x)]$
$=\sup_{(\omega,x)\in\Omega\times Y}\left|\frac{\psi\circ T_{\omega}^{-1}(x)}{{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}}\right|\int_{Y}\kappa(x,y)m(y)\mathrm{d}y$
$\leq\sup_{(\omega,x)\in\Omega\times Y}\left|\frac{\psi\circ T_{\omega}^{-1}(x)}{{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}}\right|\|\kappa\|_{\infty}\mu(Y)<\infty$
for some bounded and measurable function $\kappa:Y\times Y\to\mathbb{R}$. For a detailed construction of $\kappa$ see [BassolsCastroLamb2024, Theorem 2.17]. Hence, $m_{\varepsilon}\in L_{\infty}(Y,\mathrm{Leb}_{Y})$ and the proof is concluded by observing that the linear operator
$h\in\mathcal{F}_{b}(Y)\mapsto\mathbb{E}_{\varepsilon}\left[\frac{\psi\circ T_{\omega}^{-1}(x)}{{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}}h\circ T_{\omega}^{-1}(x)\cdot\mathbbm{1}_{Y}\circ T_{\omega}^{-1}(x)\right]$
is strong Feller (see [BassolsCastroLamb2024, Proposition 2.17]), which implies that $m$ is continuous. ∎
The following three results will allow us to apply Lemma 4.2. Theorem 4.3 provides a bound on the essential spectral radius of the deterministic transfer operator on $B^{t,s}$, Lemma 4.4, which partly relies on Theorem 3.15, ensures that the spectral perturbation arguments of [KellerLiverani1999] are applicable, and Lemma 4.5 shows that the annealed transfer operator is strong Feller.
Theorem 4.3 (see [Baladi2018-book, Theorem 5.1] or [Baladi-Tsujii2008, Theorem 1.1]).
For $r>1$, let $T:V\to M$ be a $\mathcal{C}^{r}$ diffeomorphism onto its image with a hyperbolic basic set $\Lambda\subset V$ and isolating neighbourhood $V$. For any $g\in\mathcal{C}^{r-1}(\overline{V})$ supported in $V$ and all real numbers $s,t$ such that $t-(r-1)<s<0<t$, the transfer operator $\mathcal{P}_{g}\varphi=g(\varphi\circ T)$ extends to a bounded operator on $B^{t,s}(T,V)$ and the essential spectral radius $r_{\mathrm{ess}}(\mathcal{P}_{g}|_{B^{t,s}})$ satisfies
$\begin{split}r_{\mathrm{ess}}(\mathcal{P}_{g}|_{B^{t,s}})\leq Q^{t,s}(T,g)\coloneqq\exp\sup_{\mu\in\mathrm{Erg}(\Lambda,T)}\bigg{\{}h_{\mu}(T)+\chi_{\mu}\left(\frac{g}{\det(DT|_{E^{u}})}\right)\\ +\max\{t\chi_{\mu}(DT|_{E^{s}}),|s|\chi_{\mu}(DT^{-1}|_{E^{u}})\}\bigg{\}},\end{split}$
where $\mathrm{Erg}(\Lambda,T)$ denotes the set of $T$-invariant ergodic Borel probability measures on $\Lambda$, $h_{\mu}(T)$ denotes the metric entropy of $(\mu,T)$, and $\chi_{\mu}(A)\in\mathbb{R}\cup\{-\infty\}$ denotes the largest Lyapunov exponent of a linear cocycle $A$ over $T|_{\Lambda}$, with $(\log\|A\|)^{+}\in L_{1}(\mathrm{d}\mu)$.
Lemma 4.4.
Consider two finite families of cone systems $\Theta=\{\Theta_{i}\}_{i\in I},{\Theta^{\prime}}=\{{\Theta^{\prime}}_{i}\}_{i\in I}$ satisfying ${\Theta}_{i}^{\prime}<\Theta_{i}$ for all $i\in I$. Let $t-(r-1)<s<0<t$ and let $\rho_{0}=Q^{t,s}(T,g)$ from Theorem 4.3. Fix $\rho_{0}<\tilde{\rho}<r(\mathcal{P}_{g})$. Then there exists $\varepsilon_{0}>0$ and constants $C,C_{0}>0$ such that for all $m\in\mathbb{N}$ and all $0<\varepsilon<\varepsilon_{0}$ we have the following bounds:
$\|\mathcal{P}_{\varepsilon}^{m}\|_{B^{\Theta^{\prime},t-1,s-1}}\leq CC_{0}^{m},$ (7)
$\|\mathcal{P}_{\varepsilon}^{m}\varphi\|_{B^{\Theta,t,s}}\leq C{\tilde{\rho}}^{m}\|\varphi\|_{B^{\Theta,t,s}}+CC_{0}^{m}\|\varphi\|_{B^{\Theta^{\prime},t-1,s-1}},$ (8)
$\|\mathcal{P}_{\varepsilon}\varphi-\mathcal{P}_{g}\varphi\|_{B^{\Theta^{\prime},t-1,s-1}}\leq C\varepsilon\|\varphi\|_{B^{\Theta,t,s}}.$ (9)
In other words, $\mathcal{P}_{g}$ and $\mathcal{P}_{\varepsilon}$ satisfy the same Lasota-Yorke inequality. Moreover, all the assumptions of [KellerLiverani1999, Theorem 1] (which we recall in Theorem A.1) hold for $\mathcal{B}=B^{\Theta,t,s}$, the norms $\|\cdot\|=\|\cdot\|_{B^{\Theta,t,s}}$ and $|\cdot|=\|\cdot\|_{B^{\Theta^{\prime},t-1,s-1}}$, and the operators
$P_{\varepsilon}\coloneqq\mathcal{P}_{\varepsilon}\ \text{if }\varepsilon>0\ \text{and }P_{0}\coloneqq\mathcal{P}_{g}.$
Proof.
The proof of the above inequalities is included in the proof of [Baladi2018-book, Theorem 5.22 and Remark 5.23]. From Theorem 3.15 we have that $B^{\Theta,t,s}$ compactly embeds in $B^{\Theta^{\prime},t-1,s-1}$. It is then immediately observed that $\mathcal{P}_{g}$ and $\mathcal{P}_{\varepsilon}$ satisfy the Lasota-Yorke inequality with the same constants and the conditions of Theorem A.1 are fulfilled. ∎
We shall use the bounds of Lemma 4.4 later in the proof of Theorem 2.7 for vanishing weights. The final condition to be satisfied in order to apply Lemma 4.2 is for the operator to be strong Feller.
Lemma 4.5 ([BassolsCastroLamb2024, Proposition 2.17]).
For any compact set with non-empty interior $Y\subset M$ and any non-negative continuous function $\varphi:Y\to\mathbb{R}$, the operator
$\mathcal{P}_{\varepsilon,Y}:\mathcal{F}_{b}(Y)$ $\to\mathcal{F}_{b}(Y)$
$h$ $\mapsto\varphi(x)\mathbb{E}_{x}\left[h\circ X_{1}^{\varepsilon}\cdot\mathbbm{1}_{Y}\circ X_{1}^{\varepsilon}\right],$
where $\mathcal{F}_{b}(Y)\coloneqq\{h:Y\to\mathbb{R};\ h\text{ is bounded and measurable}\}$, is strong Feller, i.e. for every $h\in\mathcal{F}_{b}(Y)$ we have that $\mathcal{P}_{\varepsilon,Y}h$ is a continuous function.
We are ready to prove the local existence and uniqueness of quasi-ergodic measures as well as conditioned stochastic stability as in Theorem 2.7 for Hölder weights vanishing on the boundary of the isolating neighbourhood $V$.
Proof of Theorem 2.6 and Theorem 2.7 for vanishing Hölder weights.
We consider two finite families of cone systems $\Theta=\{\Theta_{i}\}_{i\in I},{\Theta^{\prime}}=\{{\Theta^{\prime}}_{i}\}_{i\in I}$ satisfying ${\Theta}_{i}^{\prime}<\Theta_{i}$ for all $i\in I$. Let $s-1<s^{\prime}<s$ and $t-1<t^{\prime}<t$ such that $t-s^{\prime}<r-1$, and consider the three anisotropic scales $B^{\Theta,t,s}\subset B^{\Theta^{\prime},t^{\prime},s^{\prime}}\subset B^{\Theta^{\prime},t-1,s-1}$.
Let $\varepsilon>0$ be fixed and let $\mu_{\varepsilon}\in(B^{\Theta,t,s})^{*},g_{\varepsilon}\in B^{\Theta,t,s}$ be such that
$\mathcal{P}^{*}_{\varepsilon}\mu_{\varepsilon}=\lambda_{\varepsilon}\mu_{\varepsilon},\enskip\mathcal{P}_{\varepsilon}g_{\varepsilon}=\lambda_{\varepsilon}g_{\varepsilon},\text{ and }\mu_{\varepsilon}(g_{\varepsilon})=1,$
with $\lambda_{\varepsilon}=r(\mathcal{P}_{\varepsilon})$. For $\delta>0$ small enough such that $B_{\delta}(\lambda)\cap\sigma(\mathcal{P}_{g})=\{\lambda\}$, define the spectral projection
$\Pi_{\varepsilon}^{(\lambda,\delta)}\coloneqq\frac{1}{2\pi i}\int_{\partial B_{\delta}(\lambda)}(z-\mathcal{P}_{\varepsilon})^{-1}\mathrm{d}z.$ (10)
Observe that $\Pi_{\varepsilon}^{(\lambda,\delta)}\varphi=\mu_{\varepsilon}(\varphi)\cdot g_{\varepsilon}$ for any $\varphi\in B^{\Theta,t,s}$. From Theorem A.1 item (2), there exists $K>0$ such that for all $\varphi\in B^{\Theta,t,s}$
$\|\Pi^{(\lambda,\delta)}_{\varepsilon}\varphi\|_{B^{\Theta,t,s}}\leq K\|\Pi^{(\lambda,\delta)}_{\varepsilon}\varphi\|_{B^{\Theta^{\prime},t-1,s-1}}.$ (11)
We renormalise $g_{\varepsilon}$ such that $\|g_{\varepsilon}\|_{(B^{\Theta,t,s})^{*}}=1$ for every $\varepsilon>0$. From (10) and (11), noting that $\Pi_{\varepsilon}^{\lambda,\delta}g_{\varepsilon}=\mu_{\varepsilon}(g_{\varepsilon})g_{\varepsilon}=g_{\varepsilon}$, it follows that there exist $K,\tilde{K}>0$ such that
$1=\|g_{\varepsilon}\|_{B^{\Theta,t,s}}\leq K\|g_{\varepsilon}\|_{B^{\Theta^{\prime},t-1,s-1}}\leq K\|g_{\varepsilon}\|_{B^{\Theta^{\prime},t^{\prime},s^{\prime}}}\leq K\tilde{K}\|g_{\varepsilon}\|_{B^{\Theta,t,s}}\leq K\tilde{K},$
from which we obtain that $1/K\leq\|g_{\varepsilon}\|_{B^{t^{\prime},s^{\prime}}}\leq\tilde{K}$ for all $\varepsilon>0$ small enough.
Let $\mu\in(B^{\Theta,t,s})^{*},\gamma\in B^{\Theta,t,s}$ be such that $\mathcal{P}_{g}^{*}\mu=\lambda\mu,\mathcal{P}_{g}\gamma=\lambda\gamma$ and $\mu(\gamma)=1$, with $\lambda=r(\mathcal{P}_{g})$. We show that for every $\phi\in\mathcal{C}^{\infty}(\overline{V})$, the following limit holds:
$\lim_{\varepsilon\to 0}\mu_{\varepsilon}(\phi\cdot g_{\varepsilon})=\mu(\phi\cdot\gamma).$
Recall that $B^{\Theta,t,s}$ compactly embeds in $B^{\Theta^{\prime},t^{\prime},s^{\prime}}$ by Theorem 3.15 and thus there exists a subsequence $\{g_{\varepsilon_{n}}\}_{n\in\mathbb{N}}\subset\{g_{\varepsilon}\}_{\varepsilon>0}$ with $\varepsilon_{n}\to 0$ such that $g_{\varepsilon_{n}}\to g_{0}\neq 0$ in $B^{\Theta^{\prime},t^{\prime},s^{\prime}}$. Moreover, it follows from Lemma 4.4 combined with item (3) of Theorem A.1 (see also [Baladi2018-book, Appendix A]) that $r(\mathcal{P}_{\varepsilon})=\lambda_{\varepsilon}$ converges to $\lambda=r(\mathcal{P}_{g})$ as $\varepsilon\to 0$.
Applying Lemma 4.4 to $B^{\Theta^{\prime},t^{\prime},s^{\prime}}$, there exists $C$ such that
$\|\lambda_{\varepsilon}g_{\varepsilon}-\mathcal{P}_{g}g_{\varepsilon}\|_{B^{\Theta^{\prime},t^{\prime}-1,s^{\prime}-1}}=\|\mathcal{P}_{\varepsilon}g_{\varepsilon}-\mathcal{P}_{g}g_{\varepsilon}\|_{B^{\Theta^{\prime},t^{\prime}-1,s^{\prime}-1}}\leq C\varepsilon\|g_{\varepsilon}\|_{B^{\Theta^{\prime},t^{\prime},s^{\prime}}}\leq C\tilde{K}\varepsilon.$
Therefore, we obtain that $\mathcal{P}_{g}g_{0}=\lambda g_{0}\in B^{\Theta^{\prime},t^{\prime},s^{\prime}}$, implying that $g_{0}=\gamma$. Since this holds for any subsequence we conclude that $g_{\varepsilon}\to\gamma$ in $B^{\Theta^{\prime},t^{\prime},s^{\prime}},$ as $\varepsilon\to 0$.
From item (1) of Theorem A.1, there exists $K_{1}>0$ and $\eta\in(0,1)$ such that for all $f\in B^{\Theta,t,s}$
$\|\mu_{\varepsilon}(f)g_{\varepsilon}-\mu(f)\gamma\|_{B^{\Theta^{\prime},t-1,s-1}}\leq K_{1}\varepsilon^{\eta}\|f\|_{B^{\Theta,t,s}}.$
Therefore,
$\begin{split}|\mu_{\varepsilon}(f)-\mu(f)|\cdot\|g_{\varepsilon}\|_{B^{\Theta^{\prime},t-1,s-1}}&\leq\|\mu_{\varepsilon}(f)g_{\varepsilon}-\mu(f)\gamma\|_{B^{\Theta^{\prime},t-1,s-1}}+\mu(f)\|\gamma-g_{\varepsilon}\|_{B^{\Theta^{\prime},t-1,s-1}}\\ &\leq K_{1}\varepsilon^{\eta}\|f\|_{B^{\Theta,t,s,}}+\mu(f)\|\gamma-g_{\varepsilon}\|_{B^{\Theta^{\prime},t-1,s-1}},\end{split}$
which yields
$|\mu_{\varepsilon}(f)-\mu(f)|\frac{1}{K}\leq K_{1}\varepsilon^{\eta}\|f\|_{B^{\Theta,t,s,}}+\mu(f)\|\gamma-g_{\varepsilon}\|_{B^{\Theta^{\prime},t^{\prime},s^{\prime}}}.$
As $\varepsilon\to 0$, we obtain that $\mu_{\varepsilon}\to\mu$ in the weak- topology of $(B^{\Theta,t,s})^{*}$ and of $(B^{\Theta^{\prime},t^{\prime},s^{\prime}})^{*}$ by the same argument.
Finally, let $\phi\in\mathcal{C}^{\infty}(\overline{V})$. We know that as $\varepsilon\to 0$,
[…]
To conclude the proof of the theorem, we show that $\nu_{\varepsilon}(\mathrm{d}x)=g_{\varepsilon}(x)\mu_{\varepsilon}(\mathrm{d}x)$ is a quasi-ergodic measure of $X_{\varepsilon}^{\log\psi}$ on $\overline{V}$. The operator $\mathcal{P}_{\varepsilon}$ is strong Feller so $\mathcal{P}^{2}_{\varepsilon}$ is compact (see [BassolsCastroLamb2024, Proposition 2.17]) and therefore $\mathcal{P}_{\varepsilon}:\mathcal{C}^{0}(\overline{V})\to\mathcal{C}^{0}(\overline{V})$ is quasi-compact [Yosida2, Equation (8), Theorem  4]. We argue that $\mathcal{P}_{\varepsilon}:\mathcal{C}^{0}(\overline{V})\to\mathcal{C}^{0}(\overline{V})$ has a spectral gap. To see this, observe that since $\mathcal{P}_{\varepsilon}(\mathcal{C}^{0}(\overline{V}))\subset\mathcal{C}^{0}_{0}(\overline{V})$, we obtain that $\left.\mathcal{P}\right|_{\mathcal{C}^{0}_{0}(\overline{V})}$ is a quasi-compact operator. Using that $\mathcal{C}^{\infty}_{0}(\overline{V})$ is dense in both $\mathcal{C}^{0}_{0}(\overline{V})$ and $B^{t,s}(\overline{V})$, we obtain that $\left.\mathcal{P}_{\varepsilon}\right|_{\mathcal{C}^{0}_{0}(\overline{V})}$ has a spectral gap from Proposition A.2 and from the spectral gap of $\mathcal{P}_{\varepsilon}:B^{t,s}\to B^{t,s}$ which is a consequence of item (4) of Theorem A.1. Moreover, observe that $g_{\varepsilon}\in\mathcal{C}_{0}^{0}(M)\subset\mathcal{C}^{0}(M)$ and $\mu_{\varepsilon}\in\mathcal{M}(\overline{V})$. A posteriori, we conclude that $\mathcal{P}_{\varepsilon}:\mathcal{C}^{0}(\overline{V})\to\mathcal{C}^{0}(\overline{V})$ also has a spectral gap as desired. Finally, invoking Lemma 4.2 it follows that $\nu_{\varepsilon}(\mathrm{d}x)=g_{\varepsilon}(x)\mu_{\varepsilon}(\mathrm{d}x)$ is a quasi-ergodic measure of $X_{\varepsilon}^{\log\psi}$ on $\overline{V}.$
Remark 4.6.
In the last argument in the proof of Theorem 2.6 and Theorem 2.7 for vanishing Hölder weights we have shown that $\mathcal{P}_{\varepsilon}$ has a spectral gap in $\mathcal{C}^{0}(\overline{V})$, which is crucial to apply Lemma 4.2 and is used in the following section.
5. Local quasi-ergodic measures for non-vanishing weights
Thus far, we proved the local existence and uniqueness of quasi-ergodic measures for $\varepsilon>0$ small and, importantly, for non-negative weights vanishing on the boundary of the isolating neighbourhood $V\supset\Lambda$. In this section, we extend these results to the case of non-negative weights which do not (necessarily) vanish on $\partial V$.
Given a $\mathcal{C}^{r-1}$ Hölder weight $e^{\phi}:\overline{V}\to[0,\infty)$, consider a suitable neighbourhood $\mathcal{U}\supset\Lambda$, which we specify later in this section, and let $\varphi\in\mathcal{C}^{r-1}(\overline{V})$ satisfy
$\varphi(x)=\begin{cases}0,&x\in\partial V\\ e^{\phi(x)},&x\in\mathcal{U},\end{cases}$ (12)
with $\varphi(x)>0$ for every $x\in\overline{V}\setminus\partial V$, where $\Lambda\subset\mathcal{U}\subset\overline{\mathcal{U}}\subset V$. Observe that $\varphi$ is a weight for which Section 4 applies on $V$.
The strategy we present here consists in constructing a suitable neighbourhood $\mathcal{U}$ of $\Lambda$ satisfying $\Lambda\subset\mathcal{U}\subset\overline{\mathcal{U}}\subset V$ and such that the quasi-ergodic measure of:
  • $X_{\varepsilon}^{\log\varphi}$ killed outside of $\overline{V}$ and with weight $\log\varphi$,
  • $X_{\varepsilon}^{\phi}$ killed outside of $\overline{V}$ and with weight $\phi$, and
  • $\widetilde{X}_{\varepsilon}^{\phi}$ killed outside of $\overline{\mathcal{U}}$ and with weight $\phi$
all coincide on $\mathcal{U}$. So far, we have shown the existence, uniqueness and stochastic stability of the quasi-ergodic measure for $X_{\varepsilon}^{\log\varphi}$, which we proved is built from the left and right dominant eigenfunctions of $\mathcal{P}_{\varepsilon}$. In Section 5.2, we show that the quasi-ergodic measures for $X_{\varepsilon}^{\phi}$ and $\widetilde{X}_{\varepsilon}^{\phi}$ are built analogously with eigenfunctions that can be obtained from those of $\mathcal{P}_{\varepsilon}$.
5.1. Constructing $\mathcal{U}$
Our construction of $\mathcal{U}$ is inspired by the well-established decomposition of the dynamics into recurrent and gradient-like behaviour, i.e. “Conley’s fundamental theorem of dynamical systems” [Conley1978, Shub1987, Norton1995]. We first recall some standard definitions and results.
Definition 5.1 ($\varepsilon$-pseudo-orbit).
Given $x,y\in M$ and $\varepsilon>0$, an $\varepsilon$-pseudo-orbit from $x$ to $y$ is a sequence of points $\{x_{0}=x,x_{1},\dots,x_{n}=y\},n>0,$ such that for $d(f(x_{k}),x_{k+1})<\varepsilon$ for every $k\in\{0,\dots,n-1\}$. If there exists a $\varepsilon$-pseudo-orbit $\{x_{0},\ldots,x_{n}\},n>0,$ such that $x=x_{0}=x_{n}$ then we say that $x$ is $\varepsilon$-pseudo-periodic.
Definition 5.2 (Chain recurrence, $x\sim y$).
We say that $x\in M$ is chain-recurrent for the map $T:M\to M$ if it is $\varepsilon$-pseudo-periodic for all positive $\varepsilon>0$. We denote by $R(T)$ the set of all chain-recurrent points. Moreover, we say that $x\sim y$ if for each $\varepsilon>0$, there exists an $\varepsilon$-pseudo-orbit from $x$ to $y$ and from $y$ to $x$.
Definition 5.3 (Complete Lyapunov function [Norton1995]).
A complete Lyapunov function for the space $M$ with respect to a continuous function $T:M\to M$ is a continuous, real-valued function $\Psi:M\to\mathbb{R}$ satisfying:
  1. (1)
    $\Psi$ is strictly decreasing on orbits outside the chain recurrent set;
  2. (2)
    $\Psi(R(T))$ is a compact nowhere dense subset of $\mathbb{R}$;
  3. (3)
    if $x,y\in R(T)$, then $\Psi(x)=\Psi(y)$ if and only if $x\sim y$; that is, for any $c\in R(T),$ $\Psi^{-1}(c)$ is a chain transitive component of $R(T).$
Theorem 5.4 ([Norton1995, Theorem 4]).
Let $T$ be a continuous function on a compact metric space $M$, then there exists a complete Lyapunov function $\Psi:M\to\mathbb{R}$ for $T$.
In the following, we construct the set $\mathcal{U}$. Let $\Psi:M\to\mathbb{R}$ be the complete Lyapunov function given by Theorem 5.4, observe from Theorem 5.4 item (3) and Hypothesis HL that $\Psi(x)$ is constant for every $x\in\Lambda$. Recall that $V$ is an isolating neighbourhood of $\Lambda$. Fix $\xi>0$ small and define the open set
$U_{n}\coloneqq\{x\in{V}:\,T^{i}(x)\in V\text{ and }|\Psi\circ T^{i}(x)-\Psi(\Lambda)|<\xi,\text{ for every }-n\leq i\leq n\}$
such that $U_{1}\supset U_{2}\supset\dots\supset\Lambda$ and $\Lambda=\cap_{n\in\mathbb{N}}U_{n}$. Moreover, we define a version of stable and unstable sets of $\Lambda$ for points in $V$ by
$W^{s}_{V}(\Lambda)\coloneqq\bigcap_{n\geq 0}T^{-n}(\overline{V})\quad\text{ and }\quad W^{u}_{V}(\Lambda)\coloneqq\bigcap_{n\geq 0}T^{n}(\overline{V}),$
which are both closed subsets of $V$. For $\delta>0$, denote by $B_{\delta}(A)$ the open ball of radius $\delta$ around a set $A\subset M$.
The following result ensures that points in $U_{n}$ do not return close to $\Lambda$ if they have escaped $U_{n}$.
Lemma 5.5.
For every $n>0$, if $x\in U_{n}$ then $T(x)\not\in W^{s}_{V}(\Lambda)\setminus U_{n}$.
Proof.
Let $x\in U_{n}$ and assume for a contradiction that $T(x)\in W^{s}_{V}(\Lambda)\setminus U_{n}$. Then, for all $i>0,\Psi(T^{i}(x))>\Psi(\Lambda)$ and $\Psi(\Lambda)+\xi>\Psi(T^{-n}(x))>\dots>\Psi(T(x))>\dots>\Psi(T^{n+1}(x))>\Psi(\Lambda),$ so $T(x)\in U_{n}$. ∎
Lemma 5.5 has a stochastic analogue as follows.
Lemma 5.6.
Given $N\in\mathbb{N}$, there exist $\delta,\varepsilon>0$ small enough such that for every $x\in U_{N}$, $T_{\omega}(x)\not\in B_{\delta}(W^{s}_{V}(\Lambda))\setminus U_{N}$ for all $\omega\in\Omega_{\varepsilon}.$
Proof.
Arguing by contradiction, assume that there exist sequences of positive real numbers $\{\delta_{n}\}_{n\in\mathbb{N}}$ and $\{\varepsilon_{n}\}_{\in\mathbb{N}}$ converging to 0, and a sequence of points $\{x_{n}\}_{n\in\mathbb{N}}$ such that $x_{n}\in U_{N}$ and $T_{\omega_{n}}(x_{n})\in B_{\delta_{n}}(W^{s}_{V}(\Lambda))\setminus U_{N}$ for every $n\in\mathbb{N}$, $\omega_{n}\in\Omega_{\varepsilon_{n}}$. Let $y_{n}=T_{\omega_{n}}(x_{n})\in B_{\delta_{n}}(W^{s}_{V}(\Lambda))\setminus U_{N}$. Let $x_{n}\to x^{*}\in\overline{U}_{N}$ as $n\to\infty$, with $y_{n}\to y^{*}=T(x^{*})\in W^{s}_{V}(\Lambda)\setminus U_{N}$. If $x^{*}\in U_{N}$, then this contradicts Lemma 5.5. Hence, assume that $x^{*}\in\partial U_{N}=\overline{U}_{N}\setminus U_{N}$. Since $y^{*}\in W^{s}_{V}(\Lambda)$ we have that $\Psi(T^{i}(y^{*}))>\Psi(\Lambda)$ for all $i>0$. It follows that $\Psi(\Lambda)+\xi=\Psi(T^{-N}(x^{*}))>\Psi(T^{-N}(y^{*})>\dots>\Psi(y^{*})>\dots>\Psi(T^{N}(y^{*}))>\Psi(\Lambda),$ so $y^{*}\in U_{N}$, which is a contradiction. ∎
This Lemma 5.6 is essential to obtain a correspondence between the quasi-ergodic measures of the three processes described above and constitutes the main property we require for the set $\mathcal{U}$ (see Lemmas 5.9 and 5.10 below).
Notation 5.7.
We set $\mathcal{U}\coloneqq U_{N}$ with $N\in\mathbb{N}$ such that $\Lambda\subset\mathcal{U}\subset\overline{\mathcal{U}}\subset V,$ with $V$ being the isolating neighbourhood of $\Lambda$. Moreover, we fix $\delta>0$ and $\varepsilon_{0}>0$ to those given by Lemma 5.6 and such that $(B_{\delta}(W^{s}_{V}(\Lambda))\cap B_{\delta}(W^{u}_{V}(\Lambda)))\setminus\mathcal{U}=\emptyset$.
5.2. Correspondence of quasi-ergodic measures
Consider the three operators
$\overline{\mathcal{P}}_{\varepsilon}f\coloneqq\varphi(x)\mathbb{E}_{x}\left[f\circ X_{1}^{\varepsilon}\cdot\mathbbm{1}_{\overline{V}}\circ X^{\varepsilon}_{1}\right]$
$\mathcal{P}_{\varepsilon}f\coloneqq e^{\phi(x)}\mathbb{E}_{x}\left[f\circ X_{1}^{\varepsilon}\cdot\mathbbm{1}_{\overline{V}}\circ X^{\varepsilon}_{1}\right]$
$\mathcal{P}_{\varepsilon,\mathcal{U}}f\coloneqq e^{\phi(x)}\mathbb{E}_{x}\left[f\circ\widetilde{X}_{1}^{\varepsilon}\cdot\mathbbm{1}_{\overline{\mathcal{U}}}\circ\widetilde{X}^{\varepsilon}_{1}\right]$
where $\widetilde{X}^{\varepsilon}$ is the process $X^{\varepsilon}$ absorbed outside of $\mathcal{U}$, i.e. if $\widetilde{X}^{\varepsilon}_{n}\not\in\mathcal{U}$ for some $n\in\mathbb{N}$ then $\widetilde{X}^{\varepsilon}_{n}\in\partial$ and $\widetilde{X}^{\varepsilon}_{n+1}\in\partial$, and $\varphi$ is defined in (12). Moreover, recall that by abuse of notation, we say that $T_{\omega}(\partial)=\partial$ for all $\omega\in\Omega_{\varepsilon}$; in this way, if $T_{\omega}^{i}(x)\not\in\overline{V}$ for some $i\in\mathbb{N}$, we have that $T^{i+j}_{\omega}(x)\in\partial$ for every $j\in\mathbb{N}.$ Finally, note that in this section, the operator $\mathcal{P}_{\varepsilon}$ of Section 4 is denoted by $\overline{\mathcal{P}}_{\varepsilon}$.
We are able to control the support of the peripheral eigenfunctions of $\mathcal{P}_{\varepsilon}$ as follows.
Lemma 5.8.
For every $\varepsilon<\varepsilon_{0}$, if $\overline{\mathcal{P}}_{\varepsilon}g_{\varepsilon}=\lambda_{\varepsilon}g_{\varepsilon}$ (resp. $\mathcal{P}_{\varepsilon}g_{\varepsilon}=\lambda_{\varepsilon}g_{\varepsilon})$ and $\lambda_{\varepsilon}=r(\overline{\mathcal{P}}_{\varepsilon})$ (resp. $r({\mathcal{P}}_{\varepsilon})$) , then $\{g_{\varepsilon}>0\}\subseteq\mathcal{U}\cup B_{\delta}(W^{s}_{V}(\Lambda))$.
Proof.
We show the result only for the operator $\overline{\mathcal{P}}_{\varepsilon}$ and note that the same proof applies when considering the operator $\mathcal{P}_{\varepsilon}$. Observe that there exists $N\in\mathbb{N}$ such that $T^{N}(x)\not\in\overline{V}$ for every $x\not\in\mathcal{U}\cup B_{\delta}(W^{s}_{V}(\Lambda))$. Since $\overline{V}$ is compact, there exists $\varepsilon>0$ small enough such that $T_{\omega}^{N}(x)\not\in\overline{V}$, for every $x\in\overline{V}\setminus\left(\mathcal{U}\cup B_{\delta}(W^{s}_{V}(\Lambda))\right)$ and $\omega\in\Omega_{\varepsilon}$. Hence,
$g_{\varepsilon}(x)=\frac{1}{\lambda_{\varepsilon}^{N}}\mathbb{E}_{\varepsilon}\left[\prod_{i=0}^{N}\varphi\circ T_{\omega}^{i}(x)\cdot g_{\varepsilon}\circ T_{\omega}^{N}(x)\cdot\mathbbm{1}_{\overline{V}}\circ T^{N}_{\omega}(x)\right]=0$
and the claim follows. ∎
This allows us to obtain eigenfunctions for $\mathcal{P}_{\varepsilon,\mathcal{U}}$ from those of $\overline{\mathcal{P}}_{\varepsilon}$ and ${\mathcal{P}}_{\varepsilon}$ simply restricting them on $\mathcal{U}$.
Lemma 5.9.
For every $\varepsilon<\varepsilon_{0}$ and $g_{\varepsilon}\in\ker(\overline{\mathcal{P}}_{\varepsilon}-\lambda_{\varepsilon})$, with $\lambda_{\varepsilon}=r(\overline{\mathcal{P}}_{\varepsilon})$ we have that $g_{\varepsilon}\mathbbm{1}_{\mathcal{U}}\neq 0$ and $\mathcal{P}_{\varepsilon,\mathcal{U}}(g_{\varepsilon}\mathbbm{1}_{\mathcal{U}})=\lambda_{\varepsilon}g_{\varepsilon}\mathbbm{1}_{\mathcal{U}}$. The same statement holds for $\mathcal{P}_{\varepsilon}$ instead of $\overline{\mathcal{P}}_{\varepsilon}$.
Proof.
We show the result only for the operator $\overline{\mathcal{P}}_{\varepsilon}$ and note that the same proof applies when considering the operator $\mathcal{P}_{\varepsilon}$. Given $x\in\mathcal{U}$, from Lemma 5.6 we know that $T_{\omega}(x)\not\in B_{\delta}(W^{s}_{V}(\Lambda))\setminus\mathcal{U}$ for every $\omega\in\Omega_{\varepsilon}$, which implies that $\mathbbm{1}_{\mathcal{U}}\circ T_{\omega}(x)=\mathbbm{1}_{\mathcal{U}\cup B_{\delta}(W^{s}_{V}(\Lambda))}\circ T_{\omega}(x)$. Thus, for $x\in\mathcal{U}$, we have that
$\mathcal{P}_{\varepsilon,\mathcal{U}}(g_{\varepsilon}\mathbbm{1}_{\mathcal{U}})(x)$ $=e^{\phi(x)}\mathbb{E}_{\varepsilon}\left[g_{\varepsilon}\circ T_{\omega}(x)\cdot\mathbbm{1}_{\mathcal{U}}\circ T_{\omega}(x)\right]$
$=\varphi(x)\mathbb{E}_{\varepsilon}\left[g_{\varepsilon}\circ T_{\omega}(x)\cdot\mathbbm{1}_{\mathcal{U}\cup B_{\delta}(W^{s}_{V}(\Lambda))}\circ T_{\omega}(x)\right]$
$=\varphi(x)\mathbb{E}_{\varepsilon}\left[g_{\varepsilon}\circ T_{\omega}(x)\cdot\mathbbm{1}_{\overline{V}}\circ T_{\omega}(x)\right]=\overline{\mathcal{P}}_{\varepsilon}g_{\varepsilon}(x)=\lambda_{\varepsilon}g_{\varepsilon}\mathbbm{1}_{\mathcal{U}}(x).$
We show that $g_{\varepsilon}\mathbbm{1}_{\mathcal{U}}\neq 0.$ Fix $0<\varepsilon<\varepsilon_{0},$ and assume for a contradiction that $g_{\varepsilon}\mathbbm{1}_{\mathcal{U}}=0$. From Lemma 5.8, there exists $\delta>0$ such that $\{g_{\varepsilon}>0\}\subset\mathcal{U}\cup B_{\delta}(W^{s}_{V}(\Lambda)).$ Observe that for every $x\in B_{\delta}(W^{s}_{V}(\Lambda))$ there exists $N>0$ such that either $T^{i}_{\omega}\in\mathcal{U}$ for some $i\in\{1,\ldots,N\}$ or $T^{N}_{\omega}(x)=\partial$. From Lemma 5.6 we have that $T^{N}_{\omega}(x)\not\in\mathcal{B}_{\delta}(W^{s}_{V}(\Lambda))\setminus\mathcal{U}$ for every $x\in B_{\delta}(W^{s}_{V}(\Lambda))\setminus\mathcal{U}$ and $\omega\in\Omega_{\varepsilon}$. It follows that
$g_{\varepsilon}(x)=\frac{1}{\lambda_{\varepsilon}^{N}}\mathbb{E}_{\varepsilon}\left[e^{S_{N}\phi(x)}g_{\varepsilon}\circ T_{\omega}^{N}(x)\right]=0,$
since $g_{\varepsilon}(\partial)=0$ and $T_{\omega}(\partial)=\partial$ for every $\omega\in\Omega_{\varepsilon}.$
We may also go in the other direction and obtain eigenfunctions for $\overline{\mathcal{P}}_{\varepsilon}$ from those of $\mathcal{P}_{\varepsilon,\mathcal{U}}$.
Lemma 5.10.
For each eigenfunction $g$ satisfying $\mathcal{P}_{\varepsilon,\mathcal{U}}g=\lambda g$, $\lambda\neq 0$, we can induce an eigenfunction of $\overline{\mathcal{P}}_{\varepsilon}$ and of $\mathcal{P}_{\varepsilon}$.
Proof.
We prove the statement for $\overline{\mathcal{P}}_{\varepsilon}$ and note that the same proof for $\mathcal{P}_{\varepsilon}$ follows changing $\varphi$ for $e^{\phi}$ below. For each $x\in\overline{V},\omega\in\Omega_{\varepsilon}$ let us define $\sigma(\omega,x)\coloneqq\min\{n\geq 0:\,T^{n}_{\omega}(x)\in\mathcal{U}\cup\partial\}$, which is a uniformly bounded stopping time. Consider the function
$\overline{g}(x)\coloneqq\mathbb{E}_{\varepsilon}\left[\tilde{g}(\omega,x)\right]\coloneqq\mathbb{E}_{\varepsilon}\left[\frac{1}{\lambda^{\sigma(\omega,x)}}\prod_{i=0}^{\sigma(\omega,x)-1}\varphi\circ T^{i}_{\omega}(x)\cdot g\circ T^{\sigma(\omega,x)}_{\omega}(x)\right].$
Notice that if $x\in\mathcal{U}$, then $\sigma(\omega,x)=0$ and $\overline{g}(x)=g(x)$. We check that $\overline{\mathcal{P}}_{\varepsilon}\overline{g}=\lambda\overline{g}$. If $x\in\overline{V}\setminus\overline{\mathcal{U}}$, then $\sigma(\omega,x)\geq 1$ and $\sigma\circ\Theta(\omega,x)=\sigma(\omega,x)-1$. Using the Markov property, we obtain
$\overline{\mathcal{P}}_{\varepsilon}\overline{g}(x)$ $=\varphi(x)\mathbb{E}_{\varepsilon}^{\omega}\left[\mathbbm{1}_{\overline{V}}\circ T_{\omega}(x)\cdot\mathbb{E}^{\nu}_{\varepsilon}\left[\tilde{g}(\nu,T_{\omega}(x))\right]\right]$
$=\mathbb{E}_{\varepsilon}\left[\mathbbm{1}_{\overline{V}}\circ T_{\omega}(x)\frac{\lambda}{\lambda^{\sigma(\omega,x)}}\prod_{i=0}^{\sigma(\omega,x)-1}\varphi\circ T^{i}_{\omega}(x)\cdot g\circ T^{\sigma(\omega,x)}_{\omega}(x)\right]=\lambda\overline{g}(x),$
where we have used the super-index on the expectation symbol to denote the noise variable with respect to which the average is taken. If $x\in\overline{\mathcal{U}}$, given $\omega\in\Omega_{\varepsilon}$ there are two possible cases: either $T_{\omega}(x)\in\overline{\mathcal{U}}$ or $T_{\omega}(x)\in\overline{V}\setminus\overline{\mathcal{U}}$. Therefore,
$\overline{\mathcal{P}}_{\varepsilon}\overline{g}(x)$ $=\varphi(x)\mathbb{E}_{\varepsilon}\left[\mathbbm{1}_{\overline{\mathcal{U}}}\circ T_{\omega}(x)\cdot\overline{g}\circ T_{\omega}(x)\right]+\varphi(x)\mathbb{E}_{\varepsilon}\left[\mathbbm{1}_{\overline{V}\setminus\overline{\mathcal{U}}}\circ T_{\omega}(x)\cdot\overline{g}\circ T_{\omega}(x)\right]$
$=\mathcal{P}_{\varepsilon,\mathcal{U}}g(x)+\varphi(x)\mathbb{E}_{\varepsilon}\left[\mathbbm{1}_{\overline{V}\setminus\overline{\mathcal{U}}}\circ T_{\omega}(x)\cdot\overline{g}\circ T_{\omega}(x)\right]$
$=\lambda\bar{g}(x)+\varphi(x)\mathbb{E}_{\varepsilon}\left[\mathbbm{1}_{\overline{V}\setminus\overline{\mathcal{U}}}\circ T_{\omega}(x)\cdot\overline{g}\circ T_{\omega}(x)\right].$
To prove the claim it is sufficient to show that $\mathbb{E}_{x}\left[\mathbbm{1}_{\overline{V}\setminus\overline{\mathcal{U}}}\circ T_{\omega}(x)\cdot\overline{g}\circ T_{\omega}(x)\right]=0$ for every $x\in{\mathcal{U}}$, which is equivalent to showing that $\overline{g}\circ T_{\omega}(x)=0$ if $T_{\omega}(x)\not\in\mathcal{U}$ for each $x\in{\mathcal{U}}$. From Lemma 5.6, since $T_{\omega}(x)\not\in\mathcal{U}$ we have that $T_{\omega}(x)\not\in B_{\delta}(W^{s}_{V}(\Lambda))$ so there exists $N>0$ sufficiently large such that for any $\nu\in\Omega_{\varepsilon}$, with $\varepsilon>0$ small enough, it holds that $T^{N}_{\nu}\circ T_{\omega}(x)\in\partial$. In particular, $g\circ T^{\sigma(\nu,x)}_{\nu}(T_{\omega}(x))=0$ for all $\nu\in\Omega_{\varepsilon}$ so $\overline{g}\circ T_{\omega}(x)=0$. ∎
With Lemma 4.2 in mind, we show that $\mathcal{P}_{\varepsilon}$ and $\mathcal{P}_{\varepsilon,\mathcal{U}}$ have a spectral gap in $\mathcal{C}^{0}$ and that their spectral radii coincide with that of $\overline{\mathcal{P}}_{\varepsilon}$.
Lemma 5.11.
The operator $\mathcal{P}_{\varepsilon,\mathcal{U}}$ has a spectral gap in $\mathcal{C}^{0}(\overline{\mathcal{U}})$ and $r(\mathcal{P}_{\varepsilon,\mathcal{U}})=r(\overline{\mathcal{P}}_{\varepsilon})$.
Proof.
Observe that for all $x\in\mathcal{U}$ and $f\in\mathcal{C}^{0}(\overline{V}),$ we have $\overline{\mathcal{P}}_{\varepsilon}(f\mathbbm{1}_{\mathcal{U}})(x)\geq\mathcal{P}_{\varepsilon,\mathcal{U}}(f\mathbbm{1}_{\mathcal{U}})(x)$ so that $r(\mathcal{P}_{\varepsilon,\mathcal{U}})\leq r(\overline{\mathcal{P}}_{\varepsilon})$. On the other hand, Lemma 5.9 provides $r(\mathcal{P}_{\varepsilon,\mathcal{U}})\geq r(\overline{\mathcal{P}}_{\varepsilon})$, so $r(\mathcal{P}_{\varepsilon,\mathcal{U}})=r(\overline{\mathcal{P}}_{\varepsilon})$.
The operator $\mathcal{P}_{\varepsilon,\mathcal{U}}$ is strong Feller and thus $\mathcal{P}^{2}_{\varepsilon,\mathcal{U}}$ is compact (see [BassolsCastroLamb2024, Proposition 2.17]). Moreover, $\frac{1}{\lambda_{\varepsilon}}\overline{\mathcal{P}}_{\varepsilon}$ is power-bounded since it has a spectral gap in $\mathcal{C}^{0}(\overline{V})$ (see Remark 4.6), and so is $\frac{1}{\lambda_{\varepsilon}}\mathcal{P}_{\varepsilon,\mathcal{U}}$. Finally, since there is a $1$-to-$1$ correspondence between the eigenvalues of $\mathcal{P}_{\varepsilon,\mathcal{U}}$ and $\overline{\mathcal{P}}_{\varepsilon}$ (Lemmas 5.9 and 5.10), this provides the presence of a spectral gap in $\mathcal{C}^{0}(\overline{\mathcal{U}})$ for $\mathcal{P}_{\varepsilon,\mathcal{U}}$. ∎
Lemma 5.12.
The operator $\mathcal{P}_{\varepsilon}$ has a spectral gap in $\mathcal{C}^{0}(\overline{V})$ and $r(\mathcal{P}_{\varepsilon})=r(\mathcal{P}_{\varepsilon,\mathcal{U}})=r(\overline{\mathcal{P}}_{\varepsilon}).$
Proof.
$\mathcal{P}_{\varepsilon}$ is a strong Feller operator and so quasi-compact. Using a similar construction to that in Lemma 5.10, we may induce eigenvectors for $\mathcal{P}_{\varepsilon}$ from those of $\mathcal{P}_{\varepsilon,\mathcal{U}}$. Moreover, $|\mathcal{P}_{\varepsilon,\mathcal{U}}f|\leq|\mathcal{P}_{\varepsilon}f|$ so that $r(\mathcal{P}_{\varepsilon,\mathcal{U}})\leq r(\mathcal{P}_{\varepsilon})$, and following Lemmas 5.8 and 5.9 we have that $r(\mathcal{P}_{\varepsilon})\leq r(\mathcal{P}_{\varepsilon,\mathcal{U}})$. It follows that $r(\mathcal{P}_{\varepsilon})=r(\mathcal{P}_{\varepsilon,\mathcal{U}})$.
We show that $\mathcal{P}_{\varepsilon}$ does not have any Jordan blocks on the peripheral spectrum. Assume that $f_{\varepsilon}$ satisfies $\mathcal{P}_{\varepsilon}f_{\varepsilon}=\lambda_{\varepsilon}f_{\varepsilon}+g_{\varepsilon}$, with $g_{\varepsilon}$ an eigenfunction of $\mathcal{P}_{\varepsilon}$ associated with the eigenvalue $\lambda_{\varepsilon}=r(\mathcal{P}_{\varepsilon})$. Following the same steps as in the proof of Lemma 5.8 we have that $\mathcal{P}_{\varepsilon,\mathcal{U}}(f_{\varepsilon}\mathbbm{1}_{\mathcal{U}})=\lambda_{\varepsilon}f_{\varepsilon}\mathbbm{1}_{\mathcal{U}}+g_{\varepsilon}\mathbbm{1}_{\mathcal{U}}$. However, $g_{\varepsilon}\mathbbm{1}_{\mathcal{U}}\neq 0$ and $\mathcal{P}_{\varepsilon,\mathcal{U}}$ does not have Jordan blocks as it is power-bounded, thus reaching a contradiction. ∎
Bringing all these properties together, we conclude with the main result of this section. Corollary 5.13 provides a clear link between the quasi-ergodic measures of the three processes considered and extends the existence and uniqueness of quasi-ergodic measures for non-vanishing weights.
Corollary 5.13.
Each of the three processes:
  • $X_{\varepsilon}^{\log\varphi},$ killed outside of $\overline{V}$ and with weight $\varphi$,
  • $X_{\varepsilon}^{\phi},$ killed outside of $\overline{V}$ and with weight $e^{\phi}$, and
  • $\widetilde{X}_{\varepsilon}^{\phi},$ killed outside of $\overline{\mathcal{U}}$ and with weight $e^{\phi}$
admits the same quasi-ergodic $\nu_{\varepsilon}$ measure, which is supported on $\overline{\mathcal{U}}$.
Proof.
The transfer operators $\overline{\mathcal{P}}_{\varepsilon},\mathcal{P}_{\varepsilon},$ and $\mathcal{P}_{\varepsilon,\mathcal{U}}$ are all strong Feller (see Lemma 4.5), have a spectral gap in $\mathcal{C}^{0}$ (see Remark 4.6 and Lemmas 5.12 and 5.11) and their left and right eigenfunctions coincide in $\mathcal{U}$ (see Lemmas 5.9 and 5.10). From Lemma 5.8 we obtain that dominant eigenfunctions of $\overline{\mathcal{P}}_{\varepsilon}$ and $\mathcal{P}_{\varepsilon}$ are supported on $\overline{\mathcal{U}}\cup B_{\delta}(W^{s}_{V}(\Lambda)$ (recall the definition of $\delta$ from Notation 5.7). Analogously, we also obtain that the dominant eigenmeasures of $\overline{\mathcal{P}}_{\varepsilon}^{*}$ and $\mathcal{P}_{\varepsilon}^{*}$ are supported on $\mathcal{U}\cup B_{\delta}(W_{V}^{u}(\Lambda)).$ From Lemma 4.2 we conclude that the quasi-ergodic measures of $X_{\varepsilon}^{\log\varphi},X_{\varepsilon}^{\phi},$ and $\widetilde{X}_{\varepsilon}^{\phi}$ are supported on $\mathcal{U}$ and therefore coincide. ∎
6. Proof of the theorems
6.1. Local conditioned stochastic stability
We have developed the necessary ingredients in Sections 3 and 5 to show the existence and conditioned stochastic stability of a quasi-ergodic measure $\nu_{\varepsilon}^{\phi}$ for the process $X_{\varepsilon}^{\phi}$ conditioned upon not leaving $\overline{V}$. Recall that this measure satisfies $\nu_{\varepsilon}^{\phi}=g_{\varepsilon}(x)\mu_{\varepsilon}(\mathrm{d}x)$, where $g_{\varepsilon}\in\ker(\mathcal{P}_{\varepsilon}-\lambda_{\varepsilon})\cap\mathcal{C}^{0}_{+}(\overline{V})$ and $\mu_{\varepsilon}(\mathrm{d}x)=m_{\varepsilon}(x)\mathrm{d}x\in\ker(\mathcal{P}_{\varepsilon}^{*}-\lambda_{\varepsilon})$ with $m_{\varepsilon}\in\mathcal{C}^{0}_{+}(\overline{V})$ (see Lemma 4.2 (3)). Here, the transfer operator considered is
$\mathcal{P}_{\varepsilon}:\mathcal{C}^{0}(\overline{V})$ $\to\mathcal{C}^{0}(\overline{V}),$
$f(x)$ $\mapsto e^{\phi(x)}\mathbb{E}_{\varepsilon}\left[f\circ T_{\omega}(x)\cdot\mathbbm{1}_{\overline{V}}\circ T_{\omega}(x)\right]$
and $\lambda_{\varepsilon}=r(\mathcal{P}_{\varepsilon})$. Following from Corollary 5.13, it is only left to show that the quasi-ergodic measure $\nu_{\varepsilon}^{\phi}$ satisfies $\mathrm{supp}\,\nu_{\varepsilon}^{\phi}\supset\Lambda$ to conclude the proof of Theorem 2.6. We dedicate the rest of Section 6.1 to this statement. Recall that by abuse of notation if $T^{n}_{\omega}(x)\not\in\overline{V}$ we say that $T^{n+m}_{\omega}(x)\not\in\overline{V}$ for every $m\in\mathbb{N}.$
Proof of Theorem 2.6.
We may distinguish three cases:
  1. (i)
    $\{g_{\varepsilon}(x)>0\}$ for all $x\in\Lambda$ and $\Lambda\subset\mathrm{supp}\,\mu_{\varepsilon}$,
  2. (ii)
    $g_{\varepsilon}(x)=0$ for some $x\in\Lambda$, and
  3. (iii)
    there exists $x\in\Lambda$ such that $x\not\in\mathrm{supp}\,\mu_{\varepsilon}$.
In case (i), the result follows immediately. Case (iii) reduces to case (ii) when considering the inverse random dynamics $T_{\omega}^{-1}$ instead of $T_{\omega}$, so we only focus on (ii). The following notation is used throughout the proof:
$A_{\varepsilon}^{+}\coloneqq\overline{V}\cap\bigcup_{\omega\in\Omega_{\varepsilon}}\bigcup_{n\geq 0}T_{\omega}^{n}(\Lambda),\qquad A_{\varepsilon}^{-}\coloneqq\overline{V}\cap\bigcup_{\omega\in\Omega_{\varepsilon}}\bigcup_{n\geq 0}(T_{\omega}^{n})^{-1}(\Lambda),\qquad A_{\varepsilon}\coloneqq A_{\varepsilon}^{+}\cup A_{\varepsilon}^{-},$
and for every $\omega\in\Omega_{\varepsilon}$ we set
$\Lambda_{\omega}\coloneqq\bigcap_{n\in\mathbb{Z}}T^{n}_{\omega}(\overline{V}),\qquad W_{V}^{s}(\omega,\Lambda_{\omega})\coloneqq\bigcap_{n\geq 0}(T^{n}_{\omega})^{-1}(\overline{V}).$
Observe that $A_{\varepsilon}$ is (totally) invariant under the dynamics conditioned upon survival, i.e. for all $n\in\mathbb{Z}$
$\overline{V}\cap\bigcup_{\omega\in\Omega_{\omega}}T_{\omega}^{n}(A_{\varepsilon})\subset A_{\varepsilon}$
and so is $\overline{A_{\varepsilon}}$ since $T_{\omega}$ is a diffeomorphism for each $\omega\in\Omega_{\varepsilon}$. We divide the proof into $6$ steps to reach a contradiction. assuming $g_{\varepsilon}(x)=0$ for some $x\in\Lambda$.
Step 1.
We show that $g_{\varepsilon}(y)=0$ for all $y\in\overline{A^{+}_{\varepsilon}}$.
Proof of Step 1.
Recall that for every $z\in\overline{V}$
$g(z)=\frac{1}{\lambda_{\varepsilon}}e^{\phi(z)}\mathbb{E}_{\varepsilon}[g_{\varepsilon}\circ T_{\omega}(z)\cdot\mathbbm{1}_{\overline{V}}\circ T_{\omega}(z)].$
In this way, since $g_{\varepsilon}$ is a continuous function, we obtain that if $g_{\varepsilon}(z)=0$ for some $z\in\overline{V}$ then $g_{\varepsilon}\circ T_{\omega}(z)=0$ for every $\omega\in\Omega_{\varepsilon}$ satisfying $T_{\omega}(z)\in\overline{V}$. Repeating the previous observation we obtain that $g_{\varepsilon}\circ T_{\omega}^{n}(z)=0$ for every $\omega\in\Omega_{\varepsilon}$ such that $T_{\omega}^{n}(z)\in\overline{V}.$
Since $T$ is mixing on $\Lambda$, $x\in\Lambda$ and $g_{\varepsilon}(x)=0$, the previous paragraph implies that $\left.g_{\varepsilon}\right|_{\Lambda}\equiv 0$. Therefore, $g_{\varepsilon}$ must also vanish on $\overline{V}\cap\bigcup_{\omega\in\Omega_{\varepsilon}}\bigcup_{n\geq 0}\bigcup_{y\in\Lambda}T^{n}_{\omega}(y)=A_{\varepsilon}^{+}.$ Since $g_{\varepsilon}$ is a continuous function we obtain that $g_{\varepsilon}(y)=0$ for every $y\in\overline{A_{\varepsilon}^{+}}.$
Step 2.
There exists a non-negative function $f_{\varepsilon}\in\mathcal{C}^{0}(\overline{V})\setminus\{0\}$ such that $\mathcal{P}_{\varepsilon}f_{\varepsilon}=\sigma_{\varepsilon}f_{\varepsilon}$ for some $\sigma_{\varepsilon}\in(0,\lambda_{\varepsilon})$ and $\{f_{\varepsilon}>0\}\supset A_{\omega}^{-}.$
Proof of Step 2.
Consider the operator $\mathcal{G}_{\varepsilon}\coloneqq\mathcal{P}_{\varepsilon,\overline{A_{\varepsilon}}}:\mathcal{C}^{0}(\overline{A_{\varepsilon}})\to\mathcal{C}^{0}(\overline{A_{\varepsilon}})$. We start by establishing some properties of $\mathcal{G}_{\varepsilon}$. The choice of the random perturbations $T_{\omega}$ in Section 2.2 ensures that $\Lambda\subset\mathrm{Int}(A_{\varepsilon})$. By Lemma 4.5, the operator $\mathcal{G}_{\varepsilon}$ is a positive strong Feller operator and, applying [BassolsCastroLamb2024, Proposition 2.17], we can conclude that $\mathcal{G}_{\varepsilon}^{2}$ is compact. We claim that $\mathcal{G}_{\varepsilon}$ has a positive spectral radius. Indeed, since $T(\Lambda)=\Lambda$ and using the nature of the perturbations $T_{\omega}$ (recall Section 2.2), there exist constants $b>0$ small and $c>0$ such that
[…]
Hence, $\mathcal{G}_{\varepsilon}$ is a positive operator and, for every $z\in\Lambda$, we have that
$\|\mathcal{G}_{\varepsilon}^{n}\mathbbm{1}_{B_{b}(\Lambda)}\|_{\infty}\geq\mathcal{G}_{\varepsilon}^{n}\mathbbm{1}_{B_{b}(\Lambda)}(z)\geq c^{n}\mathbbm{1}_{B_{b}(\Lambda)}(z)=c^{n}.$
It follows that the spectral radius satisfies $r(\mathcal{G}_{\varepsilon})\geq c>0$.
Since $\mathcal{G}_{\varepsilon}^{2}$ is a compact positive operator with positive spectral, from the Krein-Rutman theorem (see [Meyer-Nieberg1991, Theorem 4.1.4]), we obtain that there exists a non-negative function $\overline{f}_{\varepsilon}\in\mathcal{C}^{0}(\overline{A_{\varepsilon}^{+}})\setminus\{0\}$ such that $\mathcal{G}_{\varepsilon}\overline{f}_{\varepsilon}=\sigma_{\varepsilon}\overline{f}_{\varepsilon}$ where $\sigma_{\varepsilon}\coloneqq r(\mathcal{G}_{\varepsilon})$. Observe that $\overline{f}_{\varepsilon}(z)>0$ for each $z\in\Lambda$, otherwise from the same argumentation provided in Step 1 we would obtain that $\overline{f}_{\varepsilon}(y)\equiv 0$ for each $y\in\overline{A_{\varepsilon}^{+}}$. Recalling that $\overline{A_{\varepsilon}^{+}}$ is $T_{\omega}$-forward invariant and repeating proof of Lemma 5.10, we can construct a non-negative continuous function $f_{\varepsilon}\in\mathcal{C}^{0}(\overline{V})$ such that $\mathcal{P}_{\varepsilon}f_{\varepsilon}=\sigma_{\varepsilon}f_{\varepsilon}$ and $\left.f_{\varepsilon}\right|_{\overline{A_{\varepsilon}^{+}}}=\overline{f}_{\varepsilon}.$ Observe that $f_{\varepsilon}\neq g_{\varepsilon}$, because $f_{\varepsilon}(z)\neq 0$ for every $z\in\Lambda$, in particular we obtain that $0<\sigma_{\varepsilon}<\lambda_{\varepsilon}$.
We now show that $f_{\varepsilon}$ in positive in $A_{\varepsilon}^{-}$. Given $x\in A_{\varepsilon}^{-}$ there exists $n\in\mathbb{N}$, $\nu\in\Omega_{\varepsilon}$ and $y\in\Lambda$ such that $x=(T_{\nu}^{n})^{-1}(y)$. Hence,
$f_{\varepsilon}(x)$ $=f_{\varepsilon}((T_{\nu}^{n})^{-1}(y))=\frac{1}{\sigma_{\varepsilon}^{n}}\mathcal{P}^{n}_{\varepsilon}f_{\varepsilon}\left((T_{\nu}^{n})^{-1}(y)\right)$
$=\frac{1}{\sigma_{\varepsilon}^{n}}\mathbb{E}_{\varepsilon}\left[e^{S_{n}\phi\circ(T_{\nu}^{n})^{-1}(y)}f_{\varepsilon}\circ T_{\omega}^{n}\circ(T_{\nu}^{n})^{-1}(y)\cdot\mathbbm{1}_{\overline{V}}\circ T_{\omega}^{n}\circ(T_{\nu}^{n})^{-1}(y)\right]>0,$
since $0<f_{\varepsilon}(y)=f_{\varepsilon}\circ T_{\nu}^{n}\circ(T_{\nu}^{n})^{-1}(y)$. ∎
Step 3.
We show that $\nu_{\varepsilon}(\overline{A_{\varepsilon}})=0$.
Proof of Step 3.
Let $f_{\varepsilon}\in\mathcal{C}^{0}(\overline{V})$ be the function constructed in Step 2. Since
$\int_{\overline{V}}f_{\varepsilon}(x)\,\mu_{\varepsilon}(\mathrm{d}x)$ $=\frac{1}{\sigma_{\varepsilon}}\int_{\overline{V}}\mathcal{P}_{\varepsilon}f_{\varepsilon}(x)\,\mu_{\varepsilon}(\mathrm{d}x)=\int_{\overline{V}}f_{\varepsilon}(x)\,\mathcal{P}_{\varepsilon}^{*}\mu_{\varepsilon}(\mathrm{d}x)=\frac{\lambda_{\varepsilon}}{\sigma_{\varepsilon}}\int_{\overline{V}}f_{\varepsilon}(x)\,\mu_{\varepsilon}(\mathrm{d}x),$
we conclude that $0=\int_{\overline{V}}f_{\varepsilon}\,\mathrm{d}\mu_{\varepsilon}=\int_{\overline{V}}f_{\varepsilon}(x)\,m_{\varepsilon}(x)\mathrm{d}x$. Since $f_{\varepsilon}$ is non-negative and not identically zero, it follows that
$\mu_{\varepsilon}\left(\{f_{\varepsilon}>0\}\right)=\int_{\{f_{\varepsilon}>0\}}\mu_{\varepsilon}(\mathrm{d}x)=0.$
From Step 2, the above equation implies that $\Lambda\cap\operatorname{supp}(\mu_{\varepsilon})=\emptyset$, and hence $\left.m_{\varepsilon}\right|_{\Lambda}=0.$
Recalling from Lemma 4.2 item (3) that
$m_{\varepsilon}(x)=\frac{1}{\lambda_{\varepsilon}^{n}}\mathbb{E}_{\varepsilon}\left[\mathbbm{1}_{Y}\circ T_{\omega}^{-1}(x)\frac{e^{\phi\circ T_{\omega}^{-1}(x)}}{|\det DT_{\omega}|\circ T_{\omega}^{-1}(x)}m_{\varepsilon}\circ T_{\omega}^{-1}(x)\right],$
we see that if $m_{\varepsilon}(y)=0$, then $m_{\varepsilon}(T_{\omega}^{-1}(y))=0$ for every $\omega\in\Omega_{\varepsilon}$ satisfying $T_{\omega}^{-1}(x)\in\overline{V}$. Iterating this argument and using the fact that $m_{\varepsilon}$ vanishes on $\Lambda$, we obtain that $m_{\varepsilon}(z)=0$ for every $z\in A_{\varepsilon}^{-}.$ Since $m_{\varepsilon}$ is continuous, it follows that $\left.m_{\varepsilon}\right|_{\overline{A_{\varepsilon}^{-}}}\equiv 0.$ On the other hand, since $g_{\varepsilon}$ vanishes on $\overline{A_{\varepsilon}^{+}}$, we conclude that
$\nu_{\varepsilon}(\overline{A_{\varepsilon}})=\nu_{\varepsilon}(\overline{A_{\varepsilon}^{+}}\cup\overline{A_{\varepsilon}^{-}})=\int_{\overline{A_{\varepsilon}^{+}}\cup\overline{A_{\varepsilon}^{-}}}g_{\varepsilon}(x)\,m_{\varepsilon}(x)\,\mathrm{d}x=0,$
yielding Step 3. ∎
Step 4.
We show that $\{g_{\varepsilon}>0\}\subset\bigcup_{\omega\in\Omega_{\varepsilon}}W^{s}_{V}(\omega,\Lambda_{\omega})$.
Proof of Step 4.
Choose $x\in\overline{V}\setminus\cup_{\omega\in\Omega_{\varepsilon}}W^{s}_{V}(\omega,\Lambda_{\omega})$ and observe that for every $\omega\in\Omega_{\varepsilon}$, there exists $n=n(\omega)>0$ finite such that $T_{\omega}^{n(\omega)}(x)\not\in\overline{V}$. By continuity of $\nu\in\Omega_{\varepsilon}\mapsto T^{n(\omega)}_{\nu}(x)\in M$, each $\omega\in\Omega_{\varepsilon}$ admits an open neighbourhood $B_{\omega}$ of $\omega$ on $\Omega_{\varepsilon}$ such that $T^{n(\omega)}_{\nu}(x)\not\in\overline{V}$ for every $\nu\in B_{\omega}$. Since $\Omega_{\varepsilon}$ is compact, there exists $N>0$ such that $T^{N}_{\omega}(x)\not\in\overline{V}$ for all $\omega\in\Omega_{\varepsilon}$, so
$g_{\varepsilon}(x)=\frac{1}{\lambda^{N}}\mathbb{E}_{\varepsilon}[e^{S_{N}\phi}g_{\varepsilon}\circ T^{N}_{\omega}(x)\cdot\mathbbm{1}_{\overline{V}}\circ T^{N}_{\omega}(x)]=0.$
Step 5.
There exists $\varepsilon>0$ small such that $W^{s}_{V}(\omega,\Lambda_{\omega})\subset A_{\varepsilon}$ for $\mathbb{P}$-almost every $\omega\in\Omega_{\varepsilon}$.
Proof of Step 5.
It follows from [Liu1998, Theorem 1.1] that there exists $\varepsilon>0$ small enough such that $\bigcup_{\omega\in\Omega_{\varepsilon}}\Lambda_{\omega}\subset V$. Let $z\in W^{s}_{V}(\omega,\Lambda_{\omega})$ for some $\omega\in\Omega_{\varepsilon}$ and observe that the following holds true:
  • $T^{n}_{\omega}(z)\in\overline{V}$ for all $n\geq 0$,
  • $\lim_{n\to\infty}\mathrm{dist}_{H}(T^{n}_{\omega}(z),\Lambda_{\theta^{n}\omega})=0$, where $\mathrm{dist}_{H}$ denotes the Hausdorff distance, and
  • there exists $\delta=\delta(\varepsilon)$ such that $\Lambda\subset B_{\delta}(\Lambda)\subset\mathrm{Int}(A_{\varepsilon})$.
Consider the open [^7] Observe that this follows from $\Lambda_{\omega}=h_{\omega}(\Lambda)$, where $\omega\mapsto h_{\omega},h_{\omega}:\Lambda\to\Lambda_{\omega}$ is a continuous map [Liu1998, Theorem 1.1]. set $B\coloneqq\{\omega\in\Omega_{\varepsilon}\,:\,\mathrm{dist}_{H}(\Lambda_{\omega},\Lambda)<\delta/2\}$, which has positive $\mathbb{P}_{\varepsilon}$-measure. It follows from Poincaré recurrence that for $\mathbb{P}_{\varepsilon}$-almost every $\omega\in\Omega_{\varepsilon}$ there exist infinitely many $n\in\mathbb{N}$ such that $\theta^{n}\omega\in B$. We may take one such $n$ large enough such that $\mathrm{dist}_{H}(T^{n}_{\omega}(z),\Lambda_{\theta^{n}\omega})<\delta/2$. Thus, $T^{n}_{\omega}(z)\in A_{\varepsilon}$ and can conclude that $z\in A_{\varepsilon}$ from the backward invariance of $A_{\varepsilon}$. ∎
Step 6.
We show that $W^{s}_{V}(\omega,\Lambda_{\omega})\subset\overline{A_{\varepsilon}}$ for all $\omega\in\Omega_{\varepsilon}$.
Proof of Step 6.
From [Liu1998, Proposition 1.3], assuming $\varepsilon>0$ is small enough, there exists an neighbourhood $V_{\underline{0}}$ of $\Lambda$ contained in $V$ and a function $H:\Omega\times V_{\underline{0}}\to M$ such that the following holds:
  1. (1)
    $\omega\in\Omega_{\varepsilon}\mapsto H(\omega,\cdot)\in\mathcal{C}^{0}(V_{\underline{0}},M)$ is continuous, where $\underline{0}=(0)_{i\in\mathbb{Z}}\in\Omega_{\varepsilon},$
  2. (2)
    for every $\omega\in\Omega$ the map $H_{\omega}(\cdot)\coloneqq H(\omega,\cdot):V_{\underline{0}}\to V_{\omega}\coloneqq H_{\omega}(V_{\underline{0}})$ is a homeomorphism with its image and $H_{\underline{0}}(\cdot)=\mathrm{Id}$, and
  3. (3)
    the diagram
[…]
    commutes. In particular $H_{\omega}(\Lambda)=\Lambda_{\omega}$ for every $\omega\in\Omega.$
In this way, we obtain that for every $\omega\in\Omega_{\varepsilon}$, $H_{\omega}(W_{V}^{s}(\Lambda)\cap V_{\underline{0}})=W_{V}^{s}(\omega,\Lambda_{\omega})\cap V_{\omega}.$
We start by showing that $W_{V}^{s}(\omega,\Lambda_{\omega})\cap V_{\omega}\subset\overline{A_{\varepsilon}}$ for each $\omega\in\Omega_{\varepsilon}$. Let $\omega\in\Omega_{\varepsilon}$. Consider a sequence $\{\omega_{n}\}_{n\in\mathbb{N}}\subset\Omega_{\varepsilon}$, where each $\omega_{n}$ satisfies $W^{s}_{V}(\omega,\Lambda_{\omega})\subset A_{\varepsilon}$, and such that $\omega_{n}\to\omega$. Observe that for each $x\in W_{V}^{s}(\omega,\Lambda_{\omega})\cap V_{\omega}$ there exists $x_{\omega}\in W_{V}^{s}(\Lambda)\cap V_{\underline{0}}$ such that $H_{\omega}(x_{\omega})=x.$ Observe that $H_{\omega_{n}}(x_{n})\in W_{V}^{s}(\omega_{n},\Lambda_{\omega_{n}})\subset A_{\varepsilon}$. Therefore $x=\lim_{n\to\infty}H_{\omega_{n}}(x_{\omega})\in\overline{A_{\varepsilon}}.$
We conclude the proof of Step 6. Recall that $\overline{A_{\varepsilon}}$ is backwards invariant and observe that
$W_{V}^{s}(\omega,\Lambda_{\omega})=\bigcup_{n\geq 0}(T_{\omega}^{n})^{-1}\left(W_{V}^{s}(\theta^{n}\omega,\Lambda_{\theta^{n}\omega})\cap V_{\theta^{n}\omega}\right)\subset\bigcup_{n\geq 0}(T_{\omega}^{n})^{-1}\overline{A_{\varepsilon}}\subset\overline{A_{\varepsilon}}$
since, as a consequence of the properties of $H_{\omega},$ there exists $\zeta>0$ such that $V_{\omega}\supset B_{\zeta}(\Lambda_{\omega})$ for every $\omega\in\Omega_{\varepsilon},$ and $T_{\omega}$ is uniformly hyperbolic on $V$. ∎
We finish the proof of the theorem by noting that
$1=\nu_{\varepsilon}(\{g_{\varepsilon}>0\})\leq\nu_{\varepsilon}\left(\bigcup_{\omega\in\Omega_{\varepsilon}}W_{V}^{s}(\omega,\Lambda\omega)\right)\leq\nu_{\varepsilon}(\overline{A_{\varepsilon}})=0,$
which is a contradiction. It follows that $g(x)>0$ and $x\in\mathrm{supp}\,\mu_{\varepsilon}$ for every $x\in\Lambda$ and so $\Lambda\subset\mathrm{supp}\,\nu_{\varepsilon}$. ∎
Proof of Theorem 2.7.
Follows from the construction of the quasi-ergodic measure of $X_{\varepsilon}^{\phi}$ on $\overline{V}$ in Theorem 2.6, which coincides with the quasi-ergodic measure of $X^{\log\psi}_{\varepsilon}$ on $V$ and the proof of Theorem 2.7 for vanishing weights in Section 4. ∎
6.2. Global conditioned stochastic stability
The results obtained so far provide a strategy for approximating equilibrium states on uniformly hyperbolic sets $\Lambda$ via quasi-ergodic measures. In turn, these may be obtained by combining the dominant eigenfunctions of the annealed transfer operator $\mathcal{P}_{\varepsilon}$ and its dual. To compute these operators, however, it is necessary to have information on the set $\Lambda$ and an isolating neighbourhood $V\supset\Lambda$ (recall the “conditioning” term $\mathbbm{1}_{V}\circ T_{\omega}(x)$ in the definition of $\mathcal{P}_{\varepsilon}$). As argued in Section 2.3, this assumption is often not valid in model-free, data-driven applications. Instead, we may only be able to identify an attracting or trapping region $\mathcal{A}$ in the state space to which trajectories converge after a sufficiently long time. In this context, the conditioning step can only be considered upon not entering this region $\mathcal{A}$.
Moreover, by construction and the results of Lemma 4.2, it follows that the quasi-ergodic limit in (1) holds for $x\in\{g_{\varepsilon}>0\}$ which is only a subset of $V$. It is only natural to ask which other points in $M$ also satisfy this limit for a given quasi-ergodic measure and how this problem can be addressed in the presence of several hyperbolic invariant sets.
We focus on these two aspects in this final section, where we build on the results obtained thus far to address the global description of a system. Here, we work under the following assumption slightly extending HG1:
Hypothesis HG2.
We say that $(T,g)$ satisfies Hypothesis HG2 if:
  1. (i)
    $(T,g)$ satisfies HG1, and
  2. (ii)
    the topological pressure on each hyperbolic basic set $\Lambda_{i}$ is different.
Let $\{\Lambda_{i}\}_{i=1}^{k}$ be the basic sets introduced at the end of Section 2.1. We shall construct a dynamical filtration of the space $M$ in the spirit of [Conley1978] and [Shub1987]. Let us recall some standard definitions.
Definition 6.1 (Adapted filtration).
A filtration adapted to $T$ is a nested family of smooth, compact, codimension 0 submanifolds $\mathbf{M}=\{M_{i}\}_{i=0}^{\ell}$, i.e. with $\emptyset=M_{0}\subset M_{1}\subset\dots\subset M_{\ell}=M,$ and such that $T(M_{i})\subset\mathrm{int}\left(M_{i}\right),i=0,\dots,\ell$.
By abuse of notation, we may think of the trapping region $\mathcal{A}$ as a subset of $M_{0}$.
For a filtration $\mathbf{M}$ adapted to $T$, we denote by $K_{i}(\mathbf{M})$ the maximal (compact) $T$-invariant subset of $M_{i}\setminus M_{i-1}$ for each $i=1,\dots,\ell$, namely:
$K_{i}(\mathbf{M})=\bigcap_{n\in\mathbb{Z}}T^{n}(M_{i}\setminus M_{i-1}).$
Definition 6.2 (Filtration ordering).
We define a preorder $\gg$ on the $\Lambda_{i}$’s as follows: $\Lambda_{i}\gg\Lambda_{j}$ if and only if $(W^{u}(\Lambda_{i})\setminus\Lambda_{i})\cap(W^{s}(\Lambda_{j})\setminus\Lambda_{j})\neq\emptyset$, i.e. there are points that are both in the unstable set if $\Lambda_{i}$ and in the stable set of $\Lambda_{j}$. If there exists a sequence $\Lambda_{i_{0}}\gg\dots\gg\Lambda_{i_{r-1}}=\Lambda_{i_{0}}$, we say that the preorder has an $r$-cycle. In the absence of cycles, we write $\Lambda_{i}>\Lambda_{j}$ if there exists a sequence such that $\Lambda_{i}\gg\dots\gg\Lambda_{j}$.
Denote by $G=(V,E)$ the graph generated by the vertices $V=\{\Lambda_{i}\}_{i=1}^{k}$ and edges $E$ given by the preorder $\gg$. To construct a total order for $\{\Lambda_{i}\}_{i=1}^{k}$, we consider the following ordering rules:
  1. (R1)
    The order inferred by $\gg$ is always preserved, i.e. if $\Lambda_{i}\gg\Lambda_{j}$ then $\Lambda_{i}>\Lambda_{j}$.
  2. (R2)
    Given two subgraphs $\mathscr{G}_{1}=(V_{1},E_{1}),\mathscr{G}_{2}=(V_{2},E_{2})\subset G$ with disjoint vertex sets, for any $\Lambda_{1}\in V_{1}$, $\Lambda_{2}\in V_{2}$, we impose that $\Lambda_{1}>\Lambda_{2}$ if
    $\max_{\Lambda\in V_{1}}P_{\mathrm{top}}(T,\Lambda,\psi)>\max_{\Lambda\in V_{2}}P_{\mathrm{top}}(T,\Lambda,\psi).$
  3. (R3)
    Within a subgraph, if there is no preorder between two basic sets $\Lambda$ and $\widetilde{\Lambda}$, we set $\Lambda>\widetilde{\Lambda}$ if $P_{\mathrm{top}}(T,\Lambda,\psi)>P_{\mathrm{top}}(T,\widetilde{\Lambda},\psi)$. This choice does not play any role in future arguments.
We apply (R1), (R2) and (R3) to the following steps:
  1. (1)
    Begin by choosing the basic set of maximal topological pressure, which we assume to be $\Lambda_{i_{0}}$ without loss of generality. Denote by $\mathscr{G}_{i_{0}}$ the graph consisting of vertices $V_{i_{0}}\coloneqq\{\Lambda_{i}\in V\,:\,\Lambda_{i}>\Lambda_{i_{0}}\}\cup\{\Lambda_{i_{0}}\}$ and their corresponding edges given by the preorder $\gg$. We apply (R1) and (R3).
  2. (2)
    Consider the graph $G_{1}=G\setminus\mathscr{G}_{i_{0}}=(V_{1},E_{1})$.
  3. (3)
    Choose the basic set in the graph $G_{1}$ of maximal topological pressure, which we denote by $\Lambda_{i_{1}}$. Denote by $\mathscr{G}_{i_{1}}$ the graph consisting of the vertices $V_{i_{1}}\coloneqq\{\Lambda_{i}\in V\setminus V_{i_{0}}\,:\,\Lambda_{i}>\Lambda_{i_{1}}\}$ and their corresponding edges. Apply (R1) and (R3) to $\mathscr{G}_{i_{1}}$ and compare it with $\mathscr{G}_{i_{0}}$ via (R2).
  4. (4)
    Consider the graph $G_{2}=G_{1}\setminus\mathscr{G}_{i_{1}}$ and repeat the previous rules on this new graph.
  5. (5)
    Repeat until $G_{n}=G_{n-1}\setminus\mathscr{G}_{i_{n-1}}=\emptyset$.
Observe that this recursion is finite as there are finitely many basic sets. Once the ordering is established, we relabel the sets $\Lambda_{i}$ such that $\Lambda_{i}>\Lambda_{j}$ if $i>j$ while also relabelling the indexes $i_{s},s\in\{0,\dots,t\},t<n$, so that they refer to the basic sets of “maximal” topological pressure identified in step (3).
Following [Shub1987], we call this a filtration ordering. Let us provide an example of the ordering proposed above.
Example 6.3.
Consider the graph $G$ in Figure 1(a), where each vertex represents a basic set $\Lambda_{i}$ and a directed edge $\Lambda_{i}\to\Lambda_{j}$ denotes that $\Lambda_{i}\gg\Lambda_{j}$. This initial node labelling denotes the ordering of $\{\Lambda_{i}\}_{i=1}^{k-1}$ by topological pressure in an ascending fashion, i.e. the node labelled 6 denotes the basic set $\Lambda_{i_{0}}$ of maximal topological pressure.
Figure 1. (a) Initial configuration, graph $G$. (b) $G_{1}=G\setminus\mathscr{G}_{i_{0}}=\mathscr{G}_{i_{1}}\sqcup\mathscr{G}_{i_{2}}$.
Following step (1), on $\mathscr{G}_{i_{0}}$ we have the order $1>4>2>7$ with $\Lambda_{i_{0}}$ the vertex labelled 7 in Figure 1(a). After removing $\mathscr{G}_{i_{0}}$ from $G$, the new vertex of maximal topological pressure corresponds to 6 in Figure 1(b), which we now denote by $\Lambda_{i_{1}}$. On $\mathscr{G}_{i_{1}}$, the order is $5>6$ and following the second rule (R2) we obtain $1>4>2>7>5>6$. The final node 3, corresponds to $\Lambda_{i_{2}}$. We have therefore built the order on $\{\Lambda_{i}\}_{i=1}^{7}$ given by $1>4>2>7>5>6>3$. Finally, we relabel the nodes so that $\Lambda_{i}>\Lambda_{j}$ if $i>j$ and use the indexes $i_{0},i_{1},i_{2}$ to denote 7, 6 and 3 from Figure 1(a), respectively. In this example, we have $t=2$, $i_{0}=4,i_{1}=2$ and $i_{2}=3$:
[…]
Recall that in Hypothesis HG1 (iii) we require that $\Lambda$ has no cycles. In particular, this assumption allows for the order construction above to be well-defined and sets us up to apply the following result.
Proposition 6.4 ([Shub1987, Theorem 2.3]).
Let $T:M\to M$ be a homeomorphism and let $\Lambda=\Lambda_{1}\sqcup\dots\sqcup\Lambda_{k}$ be the union of $k$ closed invariant sets containing all $\alpha$ and $\omega$ limit sets of $T$. Then, $\Lambda_{i}=K_{i}(\mathbf{M})$ for some filtration $\mathbf{M}$ adapted to $T$ if and only if $\Lambda$ has no cycles and the ordering by indices is a filtration ordering.
It follows that we may choose $\ell=n$ and consider a filtration $\mathbf{M}=\{M_{i}\}_{i=0}^{n}$ for which $K_{i}(\mathbf{M})=\Lambda_{i}\subset M_{i}\setminus M_{i-1}$ for every $i=1,\dots,n$.
It is immediate to see that for every $x\in W^{s}(\mathcal{A})$, there exists $\varepsilon>0$ and $N\in\mathbb{N}$ such that $T^{n}_{\omega}(x)\in\mathcal{A}$ for every $n\geq N$ and $\omega\in\Omega_{\varepsilon}$ (see Theorem 6.5 item (iii)). For such initial conditions, the transient is too short and does not hold any meaningful statistics, i.e. conditioning upon $\tau>n$ becomes the empty set for $n\geq N$ so the quasi-ergodic measure definition is ill-posed.
Let $x\in W^{s}(\Lambda_{j})\cap M_{j}$ for some $j\in\{1,\dots,n\}$ and assume that $\Lambda_{j}\in V_{i_{k}}$ for some $k\in\{0,\dots,t\}$, e.g. in the example above one could have $x\in W^{s}(\Lambda_{6})\cap M_{6}$ so $k=0$. Then, we show that the quasi-ergodic measure observed by $x$ when conditioned upon not entering any trapping region $\mathcal{A}\subset M_{i_{k}-1}$ coincides with the quasi-ergodic measure for the process starting in $M_{i_{k}}\setminus M_{i_{k}-1}$ and conditioned upon remaining there. We formalise this statement in the following theorem.
Theorem 6.5 (Global conditioned stochastic stability II, $\partial=\mathcal{A}$).
Assume that $(T,g)$ satisfies Hypothesis HG2. Let $e^{\phi}=g$ on $\Lambda$. Consider the filtration ordering presented above and let $i_{0}>\dots>i_{t}$ denote the indexes identified after the final relabelling step.
  1. (i)
    If $x\in\cup_{i_{k}\leq j<i_{k-1}}W^{s}(\Lambda_{j})\cap M_{i_{k-1}-1}$ for some $k\in\{0,\dots,t\}$, with $i_{-1}=n+1$, then for every $\varphi:M\to\mathbb{R}$ bounded and measurable,
    $\lim_{n\to\infty}\mathbb{E}_{x}^{\phi}\left[\frac{1}{n}\sum_{i=0}^{n-1}\varphi\circ X_{i}^{\phi}\,\bigg{|}\,\tau>n\right]=\int\varphi\,\mathrm{d}\nu_{\varepsilon,i_{k}}^{\phi},$
    where $\nu_{\varepsilon,i_{k}}^{\phi}$ is the unique quasi-ergodic measure of $X_{\varepsilon}^{\phi}$ on $M_{i_{k}}\setminus M_{i_{k}-1}$ when conditioned upon staying within $M_{i_{k}}\setminus M_{i_{k}-1}$ (i.e. of Theorem 2.6).
  2. (ii)
    Given $i_{k}\leq j<i_{k-1}$ and $x\in W^{s}(\Lambda_{j})$ there exists $\delta=\delta(x)$ such that for every $0<\varepsilon<\delta$ and every bounded and measurable function $\varphi:M\to\mathbb{R}$, we have that
    $\lim_{n\to\infty}\mathbb{E}_{x}^{\phi}\left[\frac{1}{n}\sum_{i=0}^{n-1}\varphi\circ X_{i}^{\phi}\,\bigg{|}\,\tau>n\right]=\int\varphi\,\mathrm{d}\nu_{\varepsilon,i_{k}}^{\phi}.$
  3. (iii)
    If $x\in W^{s}(\mathcal{A})$, there exists $\varepsilon>0$ small enough and $N\in\mathbb{N}$ such that $T^{n}_{\omega}(x)\in\mathcal{A}$ for all $n\geq N$ and all $\omega\in\Omega_{\varepsilon}$.
Observe that if $k=0$, we need only assume Hypothesis HG1 and may set $\partial$ to be any trapping region $\mathcal{A}\subset M_{i_{0}-1}$. In this case, Theorem 6.5 reduces to Theorem 2.9.
Proof of Theorem 6.5.
We prove item (i). Assume first that $x\in W^{s}(\Lambda_{j})$ for $j\geq i_{0}$ so that $k=0$ and $M_{i_{-1}-1}=M_{n}=M$. For $\varepsilon>0$ small enough, consider the global transfer operator acting on suitable functions $f:M\setminus\mathcal{A}\to\mathbb{R}$ given by
$\widehat{\mathcal{P}}_{\varepsilon}f(x)\coloneqq e^{\phi(x)}\mathbb{E}_{\varepsilon}\left[f\circ T_{\omega}(x)\cdot\mathbbm{1}_{M\setminus{\mathcal{A}}}\circ T_{\omega}(x)\right].$
Given a subset $X\subset M$ we denote by $\mathcal{P}_{\varepsilon,X}$ the transfer operator restricted to $X$, i.e.
$\mathcal{P}_{\varepsilon,X}f\coloneqq e^{\phi(x)}\mathbb{E}_{\varepsilon}\left[f\circ T_{\omega}(x)\cdot\mathbbm{1}_{X}\circ T_{\omega}(x)\right]$
for suitable $f:X\to\mathbb{R}$. Both operators are strong Feller by definition of the perturbation (cf. Section 2.2). We write $C_{i}=\overline{M_{i}\setminus M_{i-1}}$ for every $i\in\{1\dots,n\}$ and note that from Hypothesis HG2 and Lemma 5.12 we obtain that for $\varepsilon>0$ small enough the operators $\mathcal{P}_{\varepsilon,C_{i}}$ have a spectral gap in $\mathcal{C}^{0}(C_{i})$ for every $i\in\{1,\ldots,n\}$.
Observe that for $\varepsilon>0$ small enough and arguing as in Lemma 5.10, for $x\in{M_{{i_{0}}-1}}$ we have that
$\lim_{n\to\infty}\frac{1}{\lambda_{\varepsilon}^{n}}\left|\widehat{\mathcal{P}}_{\varepsilon}^{n}\mathbbm{1}_{M_{i_{0}-1}}(x)\right|$ $\leq\lim_{n\to\infty}\frac{1}{\lambda_{\varepsilon}^{n}}\sum_{i=1}^{i_{0}-1}\left|\widehat{\mathcal{P}}_{\varepsilon}^{n}\mathbbm{1}_{C_{i}}(x)\right|$
$\leq\lim_{n\to\infty}\sum_{i=1}^{i_{0}-1}\frac{K_{i}}{\lambda_{\varepsilon}^{n}}\max_{j\in\{1,\dots,i_{0}-1\}}r(\mathcal{P}_{\varepsilon,C_{j}})^{n}=0,$
for suitable constants $K_{1},\ldots,K_{i_{0}-1}>0.$ In particular, if $g_{\varepsilon,i_{0}}\in\mathcal{C}^{0}(M)$ satisfies $\widehat{\mathcal{P}}_{\varepsilon}g_{\varepsilon,i_{0}}=\lambda_{\varepsilon}g_{\varepsilon,i_{0}}$ with $\lambda_{\varepsilon}=r(\widehat{\mathcal{P}}_{\varepsilon})$, then $g_{\varepsilon,i_{0}}=\mathbbm{1}_{M\setminus M_{i_{0}-1}}g_{\varepsilon,i_{0}}+\mathbbm{1}_{M_{i_{0}-1}}g_{\varepsilon,i_{0}}=\mathbbm{1}_{M\setminus M_{i_{0}-1}}g_{\varepsilon,i_{0}}$. Since $\mathbf{M}$ is a dynamical filtration, it follows that if $x\in C_{i_{0}}$ is such that $g_{\varepsilon,i_{0}}(x)>0$ and $g_{\varepsilon,i_{0}}(T_{\omega}(x))>0$, then $T_{\omega}(x)\in C_{i_{0}}$. Therefore, $\mathcal{P}_{\varepsilon,C_{i_{0}}}g_{\varepsilon,i_{0}}=\lambda_{\varepsilon}\mathbbm{1}_{C_{i_{0}}}g_{\varepsilon,i_{0}}$. The same steps can be performed for $\mu_{\varepsilon}\in(\mathcal{C}^{0}(M))^{*}$ satisfying $\widehat{\mathcal{P}}_{\varepsilon}^{*}\mu_{\varepsilon}=\lambda_{\varepsilon}\mu_{\varepsilon}$, from which we obtain that $\mu_{\varepsilon}=\mathbbm{1}_{M_{i_{0}}}\mu_{\varepsilon}$ and ${\mathcal{P}}^{*}_{\varepsilon,C_{i_{0}}}\mu_{\varepsilon}=\lambda_{\varepsilon}\mathbbm{1}_{C_{i_{0}}}\mu_{\varepsilon}$.
An analogue of Lemma 5.10 is also applicable here: we may induce eigenfunctions of $\overline{\mathcal{P}}_{\varepsilon}$ from eigenfunctions of $\mathcal{P}_{\varepsilon,C_{i_{0}}}$, preserving their eigenvalue and setting up a link between both spectra. It follows that $\widehat{\mathcal{P}}_{\varepsilon}$ has a spectral gap in $\mathcal{C}^{0}(M)$ (cf. Lemma 5.11).
Finally, by construction of the filtration $\mathbf{M}$, for every $\varepsilon>0$ and $x\in W^{s}(\Lambda_{j}),j\geq i_{0}$, there exists $N=N(x,\varepsilon)$ such that $T^{N}_{\omega}(x)\in\Lambda_{i_{0}}$ for some $\omega\in\Omega_{\varepsilon}$. This yields $g_{\varepsilon,i_{0}}(x)>0$ and thus $\cup_{j\geq i_{0}}W^{s}(\Lambda_{j})\subseteq\{g_{\varepsilon,i_{0}}>0\}$. Applying Lemma 4.2, we conclude that for $x\in W^{s}(\Lambda_{j}),j\geq i_{0}$, the conditioned Birkhoff averages converge:
$\lim_{n\to\infty}\mathbb{E}_{x}^{\phi}\left[\frac{1}{n}\sum_{i=0}^{n-1}h\circ X_{i}^{\phi}\,\bigg{|}\,\tau>n\right]=\int h\mathrm{d}\nu_{\varepsilon,i_{0}}^{\phi},$
where $\nu_{\varepsilon,i_{0}}^{\phi}(\mathrm{d}x)=g_{\varepsilon,i_{0}}(x)\mu_{\varepsilon}(\mathrm{d}x)$ is the unique quasi-ergodic measure for the process $X^{\phi}_{\varepsilon}$ conditioned upon remaining within $M_{i_{0}}\setminus M_{i_{0}-1}$ of the local Theorem 2.7. In fact, this holds for every $x\in\{g_{\varepsilon,i_{0}}>0\}$.
For $x\in W^{s}(\Lambda_{j})$, $i_{k}\leq j<i_{k-1}$ and $k=1,\dots,t,$ we argue in a similar fashion. Let $s=i_{k-1}-1$, then the same argument above applies simply replacing $M$ by $M_{s}$ and $i_{0}$ by $i_{k}$. Observe that since $\{g_{\varepsilon,i_{k}}>0\}\subset M_{s},$ we obtain the quasi-ergodic limit for $x\in\cup_{i_{k}\geq j\geq i_{k-1}-1}W^{s}(\Lambda_{j})\cap M_{s}$.
We prove item (ii). Let $\leq j<i_{k-1}$ and $x\in W^{s}(\Lambda_{j})$. From Hypothesis HG2, for every $\varepsilon>0$ small enough there exists a filtration $\emptyset=M_{0}\subsetneq M_{1}\subsetneq M_{2}\subsetneq\ldots\subsetneq M_{n}$ of $M$ adapted to $T_{\omega}$ for every $\omega\in\Omega_{\varepsilon}$. Observe that $x\in M_{j_{0}}$ for some $j_{0}\geq j.$ From item (i), if $x\in M_{j}$ the proof is done. Assume then that $x\not\in M_{j}$ (in particular $x\not\in\Lambda_{j}$). It is sufficient to construct a new filtration $\widetilde{\mathbf{M}}$ of $M$ adapted to $T_{\omega}$ for every $\omega\in\Omega_{\delta}$ respecting the filtration ordering, i.e. $\Lambda_{i}\subset\widetilde{M}_{i}$ for every $i\in\{1,\dots,n\}$, and such that $x\in\widetilde{M}_{j},$ where $\delta=\delta(x)$ is a small constant that will be chosen appropriately during the proof.
To construct $\widetilde{\mathbf{M}}$, observe that there exist constants $N=N(x)$, $\delta=\delta(x),$ and open balls of radius $r=r(x)>0$, $B_{r}(x),B_{r}(T(x)),\ldots B_{r}(T^{N}(x)),$ such that
  • $B_{r}(T^{i}(x))\subset\mathrm{Int}(M_{k_{i}})$ for some $k_{i}$ and every $i\in\{0,1,\ldots,N\}$,
  • $B_{r}(T^{i}(x))\cap\Lambda=\emptyset$ for every $i\in\{0,1,\ldots,N\}$,
  • $B_{r}(T^{N}(x))\subset\mathrm{Int}(M_{j})$, and
  • $T^{i}_{\omega}(x)\in B_{r}(T^{i}(x))$ for every $i\in\{0,1,\ldots,N\}$ and $\omega\in\Omega_{\delta}.$
We define the new filtration by
$\widetilde{M}_{k}=M_{k}\cup\bigcup_{i=0}^{N}B_{r}(T^{i}(x))\ \text{for every }k\geq j,$
and leave $\widetilde{M}_{k}=M_{k}$ for every $k<j.$ It follows that $x\in\widetilde{M}_{j}$ and $\widetilde{\mathbf{M}}$ is a filtration of $M$ adapted to $T_{\omega}$ for every $\omega\in\Omega_{\varepsilon}$ with $0<\varepsilon<\delta$, which concludes the proof.
Finally, item (iii) follows from the observation that if $x\in W^{s}(\mathcal{A})$, there there exists $N=N(x)\in\mathbb{N}$ such that $T^{N}(x)\in\mathrm{Int}(\mathcal{A})$. By continuity we obtain that for every $\varepsilon=\varepsilon(x)>0$ small enough, depending on $x$, $T_{\omega}^{N}(x)\in\mathcal{A}$ for every $\omega\in\Omega_{\varepsilon}.$
Acknowledgements
We sincerely thank Jeroen S.W. Lamb for very valuable comments and insightful discussions throughout the preparation of this manuscript. The authors gratefully acknowledge support from the EPSRC Centre for Doctoral Training in Mathematics of Random Systems: Analysis, Modelling and Simulation (EP/S023925/1). MC’s research has been supported by an Imperial College President’s PhD scholarship and the Australian Research Council Laureate Fellowship (FL230100088).
[…]
Appendix A Some useful results from functional analysis
For the sake of completeness and the reader’s convenience, we recall two well-known spectral theorems which we use in Section 4. The first one is a particular version of the well-established perturbation theory developed in [KellerLiverani1999]. The second is a simple abstract result that provides conditions under which an operator defined on two different spaces shares the same eigenvectors for eigenvalues which are not in the essential spectrum (see also [Baladi2018-book, Appendix A.2]).
Theorem A.1 ([KellerLiverani1999, Corollary 1]).
[…]
  1. (i)
    There are constants $C_{1},M>0$ such that for every $\varepsilon\geq 0$ small enough,
    $|P_{\varepsilon}^{n}|\leq C_{1}M^{n},$
    for all $n\in\mathbb{N}$.
  2. (ii)
    There exist $C_{2},C_{3}>0$ and $\alpha\in(0,1)$, $\alpha<M,$ such that for every $\varepsilon\geq 0$
    $\|P_{\varepsilon}^{n}f\|\leq C_{2}\alpha^{n}\|f\|+C_{3}M^{n}|f|$
    for every $n\in\mathbb{N}$ and for each $f\in\mathcal{B}$.
  3. (iii)
    The closed unit ball of $(\mathcal{B},\|\cdot\|)$ is ||$\cdot$|| ⋅ |-compact in the completion of $\mathcal{B}$ with respect to the norm ||$\cdot$|| ⋅ |.
  4. (iv)
    There exists a monotone upper-semicontinuous function $\tau:[0,\infty)\to[0,\infty),$ such that $\tau(\varepsilon)>0$ if $\varepsilon>0$ and
    $\sup\{\left|P_{0}f-P_{\varepsilon}f\right|\colon f\in\mathcal{B},\|f\|\leq 1\}\leq\tau(\varepsilon)\xrightarrow[]{\varepsilon\to 0}0.$
Moreover, let $\lambda$ be an isolated simple eigenvalue of $P_{0}$ with $\lambda>\alpha$ and let $\delta>0$ be such that $B_{\delta}(\lambda)\cap\sigma(P_{0})=\{\lambda\}$. Consider the projection
$\Pi_{\varepsilon}^{(\lambda,\delta)}\coloneqq\frac{1}{2\pi i}\int_{B_{\delta}(\lambda)}(z-P_{\varepsilon})^{-1}\mathrm{d}z.$
Then, the following holds:
  1. (1)
    there exist constants $K_{1}=K_{1}(\delta,r)>0$ and $\eta>0$, such that
    $\left|\Pi_{\varepsilon}^{(\lambda,\delta)}f-\Pi_{0}^{(\lambda,\delta)}f\right|\leq K_{1}\tau(\varepsilon)^{\eta}\|f\|\ \text{for all }f\in\mathcal{B}.$
  2. (2)
    There are constants $K_{2}=K_{2}(\delta,r)>0$ and $\delta=\delta(r)>0$ such $\|\Pi_{\varepsilon}^{(\lambda,\delta)}f\|\leq K_{2}|\Pi_{\varepsilon}^{(\lambda,\delta)}|$ for all $f\in\mathcal{B}$, $\delta\in(0,\delta_{0}]$ $\varepsilon$ and small enough.
  3. (3)
    For each $\varepsilon>0$ small enough $P_{\varepsilon}$ has a unique eigenvalue $\lambda_{\varepsilon}\in B_{\delta}(\lambda)\cap\sigma(P_{\varepsilon})$, moreover $\lambda_{\varepsilon}\xrightarrow[]{\varepsilon\to 0}\lambda$, and $\lambda_{\varepsilon}$ is a simple eigenvalue of $P_{\varepsilon}$.
  4. (4)
    If $P_{0}$ has a spectral gap then each $P_{\varepsilon}$ also has a spectral gap for $\varepsilon>0$ small enough.
Proposition A.2 ([Baladi-Tsujii2008, Appendix A]).
Let $\mathcal{B}$ be a Hausdorff topological linear space and let $(\mathcal{B}_{1},\|\cdot\|_{1})$ and $(\mathcal{B}_{2},\|\cdot\|_{2})$ be Banach spaces that are continuously embedded in $\mathcal{B}$. Suppose that there is a subspace $\mathcal{B}_{0}\subset\mathcal{B}_{1}\cap\mathcal{B}_{2}$ that is dense both in the Banach spaces $(\mathcal{B}_{1},\|\cdot\|_{1})$ and $(\mathcal{B}_{2},\|\cdot\|_{2})$. Let $\mathcal{L}:\mathcal{B}\rightarrow\mathcal{B}$ be a continuous linear map, which preserves the subspaces $\mathcal{B}_{0}$, $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$. Suppose that the restrictions of $\mathcal{L}$ to $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ are bounded operators whose essential spectral radii are both strictly smaller than some number $\rho>0$. Then the eigenvalues of $\mathcal{L}|_{\mathcal{B}_{1}}$ and $\mathcal{L}|_{\mathcal{B}_{2}}$ in $\{z\in\mathbb{C}\mid|z|>\rho\}$ coincide. Furthermore, the corresponding generalised eigenspaces coincide and are contained in $\mathcal{B}_{1}\cap\mathcal{B}_{2}$.